A Complete Bibliography of Journal of the Royal Statistical Society, Series D (The Statistician):

1970–1979

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

13 October 2017 Version 1.01

Title word cross-reference

2 × 2 [373]. B [219]. β [361]. χ^{2} [499]. d′ [361]. F [499].

1 [28]. 11 [459]. 1900 [292]. 1962 [234]. 1970 [213].

21st [55]. 22nd [143]. 2nd [377, 468, 328, 129, 466].

3rd [379, 325].

4 [127]. 4th [29].

Abraham [332]. Accompanied [527]. Accountants [222]. Accounting [521, 217, 221]. Accounts [250]. Ackoff [527, 527]. Activity [83, 164]. Additives [105]. Address [404]. Adjusting [459]. Adjustment [456, 458]. Adler [117]. Administration [447]. Administrative [120]. Advanced [379, 95, 28]. Advances [304, 274]. Advisory [213]. Age [356, 479, 5]. Age-Specific [356]. Age-Stratified [479]. Aid [69, 6, 385]. Aids [339]. Airport [104, 103]. Aitchison [95]. Alan [379]. Albaum [210]. Alfredo [522]. Algebraic [84]. Algorithm [438]. Algorithms [2, 449, 3, 502, 319]. Alternatives [58]. Amendment [295].

Among [201, 175]. Analysing [430]. Analysis [376, 379, 390, 389, 417, 474, 512, 454, 155, 425, 286, 1, 39, 251, 190, 218, 506, 85, 385, 164, 287, 271, 19, 235, 254, 378, 241, 230, 354, 133, 364, 197, 217, 347, 313, 189, 427, 438, 86, 59, 191, 393, 538, 159, 455, 107, 467, 288, 392, 131, 374, 414, 420, 249, 204, 491, 315, 302, 358, 10, 108, 150, 8, 92]. Analytical [147, 471]. Ancient [330]. Anderson [235, 378]. Andrews [304]. Ang [522]. Animal [178]. Announcement [394, 402, 457]. Announcements [97, 422]. Annual [14, 55, 143]. Aoki [93]. Application [251, 306, 19, 516, 86]. Applications [247, 515, 303, 42, 444, 322, 507, 392, 273, 90]. Applied [360, 198, 74, 479, 45, 364, 44]. Appointment [82]. Appraisal [178]. Approach [417, 36, 27, 514, 530, 228, 506, 271, 378, 12, 463, 464]. Approximation [518]. Approximations [499]. Aram [42]. Arc [518]. Area [64]. Areas [229]. Argument [95]. ARIMA [459]. ARMA [496]. Armed [437]. Art [527, 435]. Arthur [473]. Aspects [218, 287, 110]. Assessing [207]. Assessment [342, 105]. Asset [219]. Assistance [448]. Associate [293]. Assumptions [347, 313]. Athletes [264]. Atkinson [520]. Auditing [531, 532, 223]. Auditor [531]. Authority [146, 165]. Automatic [511]. Average [301, 317].

B [74, 29, 520, 209, 94, 95, 127, 57, 133, 88, 333, 56]. Bachi [27]. Back [16, 32, 48, 61, 80, 100, 118, 137, 160, 171, 179, 193, 202, 214, 225, 243, 253, 260, 267, 296, 307, 335, 348, 365, 382, 396, 403, 411, 419, 434, 451, 462, 489, 498, 517, 540]. Badia [59]. Bajpai [501]. Balanced [293]. Bancroft [116]. Bandit [437]. Barrie [320, 468]. Barry [473]. Bartholomew [331]. Based [1]. Basic [522, 95, 208, 127, 148, 486, 444]. Basis [84]. Baum [359]. Bayes [239]. Bayesian [198, 299, 78]. Bearing [146]. Beginners [516]. Behaviour [425, 88]. Behavioural [472, 109, 420]. Benjamin [447]. Berelson [131]. Bernard [447, 131]. Between [496, 264]. Bibliography [23]. Bickel [304]. Binary [385, 92]. Biological [389, 83, 241, 110]. Biologists [277]. Biology [151]. Biomathematics [270]. Biomedical [135]. BIOMETRIKA [242]. Biostatistics [116, 326]. Blackith [212]. Board [64]. Bogue [24]. Bolch [209]. Boldrini [231]. Book [320, 321, 358, 376, 377, 379, 391, 474, 213, 26, 135, 23, 90, 360, 89, 231, 324, 325, 330, 74, 155, 78, 24, 291, 73, 91, 473, 526, 468, 521, 522, 22, 43, 27, 28, 29, 75, 77, 79, 136, 326, 440, 412, 465, 328, 502, 331, 210, 130, 46, 450, 506, 5, 11, 421, 520, 209, 94, 95, 211, 42, 126, 127, 157, 237, 58, 306, 115, 441, 445, 447, 449, 472, 527, 72, 116, 232, 271, 272, 276, 290, 153, 235, 305, 323, 363, 448, 503, 7, 10, 45, 378, 234]. Book [270, 241, 475, 12, 13, 57, 128, 129, 133, 148, 364, 486, 488, 292, 212, 240, 156, 238, 319, 87, 44, 132, 523, 60, 117, 92, 111, 233, 469, 332, 444, 471, 59, 76, 110, 114, 124, 149, 158, 274, 359, 470, 362, 108, 466, 504, 505, 501, 88, 294, 333, 327, 151, 322, 361, 463, 464, 393, 112, 275, 113, 152, 239, 329, 538, 109, 442, 485, 507, 508, 154, 159, 242, 277, 293, 380, 304, 467, 268, 519, 150, 269, 93, 392, 131, 476, 443, 303, 236, 446, 525, 134, 537, 420, 524, 8, 6, 9]. Book [25, 56, 375, 273]. Bookwork [132]. Boot [133]. Boots [473]. Bostwick [238]. Both [493]. Box [474, 378, 390, 400, 417, 228]. Box-Jenkins [378]. Brass [110]. Bray [500, 487]. Breiman [42]. Brian [519, 476, 524]. Brinley [469]. Britain [306, 51, 306]. British [292]. Bruyn [13]. Buckland [76]. Building [368]. Built [347, 313]. Burden [156]. Burford [12, 148]. Burington [129]. Business [41, 136, 120, 209, 271, 272, 148, 156, 345, 136, 59]. Buying [88]. By-Census [386].

C [376, 226, 95, 42, 127, 58, 306, 153, 13, 129, 133, 156, 523, 108, 504, 501, 333, 322, 277, 268]. Calculation [178, 481, 542, 541]. Calculators [207]. Calus [501]. Cameron [154]. Campbell [277]. Cancer [332]. Capital [219]. Capture [368]. Capture-Recapture [368]. Car [229]. Carl [412]. Carnival [508]. Case [298, 493, 399, 386]. Casebook [359]. Cases [406, 117]. Catastrophe [507]. Catchment [481, 542, 541]. Cattle [418]. Causal [425]. Census [54, 51, 52, 386]. Census-Taking [54]. Censuses [53]. Central [149]. Certain [374]. Chain [40]. Champernowne [11]. Change [401, 429, 357]. Changes [511]. Characterization [333]. Charles [238]. Chartered [222]. Charts [478, 27]. Chatfield [376, 504]. Chemical [83]. Cheshire [306]. Christofides [108]. City [413]. Civil [346, 40]. Clark [364]. Clarke [56]. Class [293, 40]. Classes [461]. Classical [78, 84]. Classification [197]. Clatworthy [293]. Claudia [412]. Clinical [512, 289, 515, 437]. Clustering [513, 319]. Coût [60]. Cochrane [274]. Coefficients [298]. Cold [178]. Colin [273]. Collected [275]. Collection [425, 380, 414]. College [346]. Coloured [390]. Combinatorial [107, 449]. Comment [71, 433]. Comments [389, 285, 497, 515, 401, 288, 541, 257, 316]. Common [427]. Communication [131, 122, 236]. Comparative [433]. Comparison [68, 372, 228, 310, 429, 201]. Comparisons [175]. Component [85]. Components [86]. Compounds [83]. Computation [157, 236]. Computational [393]. Computer [412, 5, 12, 25, 151]. Computers [7]. Concepts [522, 352, 42]. Conceptual [271]. Concerning [425]. Conditions [21]. Conference [213, 29]. Conflict [141]. Conjecture [435]. Connor [276, 505]. Conover [211]. Constraints [329]. Construction [3]. Constructions [107]. Consultancy [146]. Consultant [492]. Consulting [355]. Consumer [263, 294]. Contacts [166]. Containing [221]. Contemporary [463, 464]. Content [131]. Contiguity [208, 303]. Contimart [140]. Contingency [373, 387, 428, 315]. Continuous [75, 10, 269]. Contribution [199]. Control [156, 132, 530, 177, 474, 468]. Controlled [437]. Controversy [134]. Cooley [25]. Cooper [57]. Cormack [95]. Correct [418]. Correction [347]. Corrections [317, 318]. Correlate [83]. Correlation [454, 73]. Correlations [410]. Cost [535, 104]. Costing [475]. Council [448, 14]. Countries [448]. Counts [135, 1, 77, 136]. Course [22, 116, 469]. Coursework [342]. Cowden [209]. Cox [537, 133, 92]. Crabtree [156]. Craddock [5]. Cressie [497]. Criteria [514]. Criterion [374]. Critical [229, 196]. Criticism [123]. Croft [44]. Cross [538]. Cross-Tabulated [538]. Croxton [209]. Cumulative [13]. Curve [295, 265]. Cusum [478]. Cyclic [430, 154].

D

[321, 26, 89, 324, 155, 91, 521, 29, 331, 11, 209, 378, 129, 92, 233, 108, 361, 304, 150, 134, 537]. Data [390, 512, 169, 425, 286, 1, 141, 497, 430, 431, 165, 255, 85, 287, 147, 230, 354, 168, 312, 252, 380, 491, 92, 538, 476]. David [116, 234, 44, 132, 88, 112, 154]. Davies [469, 151]. Davis [87]. Day [273]. Decision [198, 506, 94, 466]. Decision-Making [506]. Decisions [89, 128, 133]. Defence [432, 518]. Definitions [95]. DeGroot [128]. Delinquent [425]. Demand [66]. Demographic [283, 250]. Demography [316, 24, 110]. Dempster [10]. Dennis [238]. Deo [449]. Departures [347, 313]. Derived [85]. Description [3]. Descriptive [464]. Design [379, 522, 510, 159, 107, 192, 386]. Designs [154, 293]. Desmond [132].

Detection [361, 511]. Determinants [494]. Determination [21]. Development [106, 220, 447]. Developments [427]. Dhrymes [363]. Diagrams [27]. Dichotomisation [142]. Dickinson [211]. Dictionary [23, 76]. Differences [306]. Differential [264]. Digital [7]. Dimensional [387]. Dimensions [429]. Dimorphism [204]. d’Indices [60]. Discrete [23, 229, 230, 75]. Discriminating [496]. Discriminatory [19]. Discussion [342]. Disney [56]. Distributed [363]. Distribution [108]. Distributional [410]. Distributions [75, 95, 23, 465, 269, 269]. Diversions [442]. Divinity [266]. Dixon [8]. Dobben [13]. Documents [60]. Donald [24, 271]. Double [298]. Dowding [127]. Dr [495]. Dr. [226]. Duckworth [6]. due [60]. Dunn [364]. During [54].

Econometric [124, 112, 328]. Econometrica [152]. Econometrics [152, 328, 113]. Economic [459, 448]. Economics [115, 254, 469, 59, 360, 11, 272, 148, 233, 519, 524]. Economists [472, 237]. Ed [328, 325, 129]. Edged [262]. Edition [446]. Editor [98, 484]. Editorial [395, 18, 35, 50, 63, 279, 453, 183]. edn [377, 468, 466, 379]. Education [349, 337, 15, 30, 480, 222]. Educational [475, 479]. Edward [449, 238]. Edwards [291, 362]. Effects [106, 493]. Eggleston [525]. Eigenvector [190]. Eilon [108]. Élaboration [60]. Elections [86]. Electricity [65]. Electronic [207]. Element [176]. Elementary [440, 59]. Elements [94, 10]. Emery [236]. Empirical [494, 239]. Employment [300]. Engineering [522, 19, 45]. Engineers [445, 472, 501, 56, 45]. Enquiries [41]. Environment [103]. Environmental [102]. Errata [201]. Erricker [95]. Error [416, 221]. Estimates [304]. Estimating [356, 37, 401]. Estimation [11, 305, 363, 219]. Estimator [357]. Evaluation [220, 219, 53]. Evanston [238]. Event [176]. Everitt [476]. Evidence [494, 525]. Examination [125, 180, 369]. Examinations [4, 31, 96, 47]. Examiner [4]. Examiners [31]. Examining [343]. Examples [359]. Exercises [25]. Experience [120, 406]. Experiences [415]. Experiment [21]. Experimental [510, 107, 192, 87]. Experimentalists [57]. Experiments [368, 140, 159, 107]. Expert [406, 405]. Exponential [497, 431, 95, 312]. Expressing [258, 318]. Extension [298]. Extensions [496].

F [90, 326, 412, 520, 209, 116, 362, 88, 154, 304, 268]. Fables [527]. Factor [251, 427, 438, 387, 155]. Factors [428]. Facts [149]. Failure [314]. Fairley [501]. Familiar [513]. FCA [500]. Federer [380]. Feigl [116]. Fiducial [311]. Fielden [475]. Financial [391, 234, 414]. Finney [26, 150]. First [309, 441, 116, 354, 14]. FIS [500]. Fisher [275, 158, 275, 510]. Fit [432, 314]. Fitting [295, 265]. Fleiss [327]. Fletcher [115]. Flows [39]. Foad [469]. Focus [149]. Food [105]. Forecasting [417, 474, 229, 459, 300, 66, 378, 256, 68, 38, 228, 310, 65, 64]. Forecasts [258, 318]. Forest [213]. Forestry [146]. Foreword [184]. Formulae [95]. Formulation [363]. Fortran [127, 237, 273]. Fourastie [60]. Fox [390, 72]. Francis [521]. Frank [487, 8, 500, 192]. Freeman [326, 361]. Freund [240]. Front [17, 33, 34, 49, 62, 81, 101, 138, 145, 161, 172, 182, 194, 215, 216, 244, 246, 261, 278, 280, 308, 336, 344, 351, 371, 384, 398, 408, 424, 426, 452, 477, 482, 509, 528, 534]. Frontiers [233]. Function [37]. Functions [300]. Further [442].

G [320, 135, 23, 90, 73, 468, 28, 77, 136, 200, 46, 11, 126, 157, 153, 323, 45, 133, 156, 238, 76, 114, 88, 151, 538, 303, 236]. Gambling [406]. Games [339]. Gandy [108, 541]. Gardner [442, 508]. Gas [20, 64]. Gauss [438]. Geary [208]. Gene [440]. General [410, 14, 55, 143, 166, 163, 95]. Generation [70, 126]. Genetics [270]. Geo [136]. Geographic [247]. Geographical [252]. Geography [285, 251, 248, 282, 288, 245, 393]. Geophysical [185]. Geophysics [186]. George [474, 303, 226]. Gerald [210]. Getis [473]. Gibbons [211]. Gillies [324]. Goal [503]. Goodness [432, 314]. Graham [538]. Graphical [27, 230, 27, 476]. Graphs [27, 436]. Grawoig [238]. Great [306]. Grocery [140]. Group [213]. Grouped [305]. Growth [234, 429]. Guenther [42]. Guise [513]. Guttman [78, 502]. Gwilym [474].

H [522, 43, 326, 328, 445, 472, 271, 276, 7, 128, 292, 505, 113, 154, 242, 293, 304]. Haber [59]. Hahn [45]. Haines [519, 524]. Haitovsky [305]. Halsey [292]. Hampei [304]. Handbook [294, 129]. Harold [325, 9]. Harrison [520]. Harshbarger [466]. Hartigan [319]. Hartley [242]. Health [166, 169, 165, 167, 481, 542, 168, 541, 170]. Henkel [134]. Henry [392]. Henschke [412]. Hertz [29]. Higher [337]. Highly [221]. Hilary [503]. Hinkley [537]. Hiorns [95]. Historical [195]. History [114, 470]. Hogg [443]. Hong [263, 386]. Hooper [152]. Hospital [164]. Housing [53]. Huang [112]. Huber [304]. Human [251, 204]. Hydrometeorology [187]. Hypotheses [425].

Ian [507]. Ichi [270]. II [135, 306, 75, 95, 470]. III [136]. IMA [496]. Implications [374]. Importance [418]. In-Built [347, 313]. Incident [1]. Independent [188]. Index [263].

Industrial [198, 492, 448]. Industries [530]. Industry [195]. Inference [325, 330, 322, 211, 321, 79, 441, 443, 9]. Information [99, 119, 181, 224, 259, 334, 381, 423, 490, 539, 162, 163, 167, 170]. Input [392]. Input-Output [392]. Inspection [468]. Inspired [192]. Institute [349, 18, 82, 96, 55, 47, 143, 125, 180]. Institutions [234]. International [29, 439]. Interpretation [400, 389, 287, 516, 380]. Interregional [306]. Interviews [415]. Intriligator [233]. Introduction [360, 74, 162, 178, 116, 217, 195, 245, 524, 26, 526, 326, 412, 519, 93, 8, 9, 25]. Introductory [469, 456, 43, 466]. Intuitive [514]. Investigating [425]. Investigation [85, 153]. Investment [178, 217]. IPC [448]. Ipsen [116]. Irving [332]. Issues [516]. IV [467].

J [26, 74, 24, 43, 29, 328, 502, 331, 5, 520, 209, 95, 211, 42, 127, 237, 276, 363, 45, 475, 133, 238, 319, 87, 132, 60, 505, 501, 327, 113, 152, 239, 538, 485, 154, 304, 268, 150, 8, 9]. Jack [370]. Jackknife [262]. James [326]. Jardine [212]. Jean [211, 364]. Jeffreys [325]. Jenkins [378, 390, 400, 417, 474, 200, 228]. Jewish [330]. Johannes [116]. John [95, 115, 156, 469, 154, 159]. Johnson [75, 269, 495]. Johnston [328]. Josas [213]. Joseph [446]. Joshi [23]. Journal [157]. Jouy [213]. Jouy-en-Josas [213]. June [96, 180, 369]. Jurg [449].

Kagan [333]. Kalton [153]. KCSG [500]. Ken [232, 290, 270]. Ken-Ichi [270]. Kendall [379, 28, 76, 470, 358, 377, 73, 323, 114]. Kessler [332]. Keys [83]. Kingdom [520, 124]. Kingman [268]. KM [500]. Kmietowicz [391]. Kojima [270]. Kong [263, 386]. Kossack [412]. Kotz [75, 269]. Krishnaiah [467]. Krutchkoff [157]. Kuh [360].

L [360, 330, 75, 79, 126, 527, 272, 276, 503, 12, 148, 238, 470, 505, 327, 109, 269, 56]. Laboratory [140]. Labrador [390]. Lady [421]. Lags [363]. Language [95]. Large [120, 516]. Larson [9]. Last [366]. Lattice [286]. Law [407, 516, 409]. Lawley [155]. Leading [264, 117]. Learning [483]. Least [438]. Lecture [238]. Lee [271]. Leo [42]. Leslie [156]. L’Esperance [272]. Letter [484]. Levin [332]. Lewis [72, 488]. Li [523]. Liddell [433]. Life [359, 446]. Lighthill [485]. Likelihood [460, 362]. Lilienfield [332]. Lindgren [94]. Lindley [89]. Linear [455, 375, 322]. Linked [169]. Linnik [333, 471]. List [96, 47, 125, 180]. Lists [120]. Literature [330, 115]. Livingstone [156]. Local [106, 165]. Location [304]. Lockwood [27]. Logistic [295, 265]. Lognormal [372]. Logrank [515]. Lohnes [25]. London [103]. Long [69, 429]. Look [417, 441]. Ltd [448]. Luck [421]. Lucy [237].

M [474, 213, 231, 73, 28, 79, 200, 210, 450, 5, 95, 449, 323, 128, 156, 233, 332, 114, 359, 501, 294, 333, 393, 152, 154, 159]. Macro [360]. Macro-Economics [360]. Made [359]. Making [506, 89]. Management [69, 472, 174, 289, 44, 108, 156]. Manager [173, 232]. Managerial [133]. Managers [290]. Managing [72]. MANOVA [514]. Manpower [36, 39, 479]. Mansfield [59]. Map [466]. Maps [27, 488]. Marascuilo [109]. Maritz [239]. Mark [241, 389]. Market [121, 117, 294]. Marketing [210, 140, 87]. Markov [285, 248, 284, 40]. Martin [442, 508]. Masanao [93]. Mask [478]. Massey [8]. Massy [88]. Master [370]. Matched [203]. Materials [238]. Mathai [450]. Mathematical [270, 156, 108, 333, 442, 212, 391, 465, 7, 240, 508]. Mathematicians [26, 472]. Mathematics [290, 350, 156, 485]. Mather [393]. Matter [16, 32, 17, 33, 48, 61, 80, 100, 34, 49, 62, 81, 118, 137, 160, 171, 101, 138, 145, 161, 179, 193, 202, 214, 172, 182, 194, 215, 225, 243, 253, 216, 244, 246, 260, 267, 296, 307, 261, 278, 280, 308, 335, 348, 365, 382, 336, 344, 351, 371, 396, 403, 411, 419, 384, 398, 408, 424, 434, 451, 462, 426, 452, 477, 489, 498, 517, 540, 482, 509, 528, 534]. Matthews [95]. Maunder [520]. Maurice [358, 377, 379, 76, 470]. Max [117]. Maxwell [155, 486, 420]. May [129]. McNicol [361]. Means [441, 201, 175]. Measure [526]. Measurement [91]. Measures [303]. Medical [162, 163, 167, 486, 511]. Medicine [91]. Medieval [330]. Meeting [14, 55, 143]. Melese [29]. Member [448]. Men [264]. Menace [350]. Meteorology [190]. Method [521, 459, 302, 231, 155]. Methodology [121, 263]. Methods [320, 198, 70, 529, 37, 153, 7, 230, 84, 158, 501, 38, 327, 252, 393, 195, 109, 177, 73, 328, 238, 112, 239]. Michael [95]. Migration [306]. Miloslav [444]. Milton [58]. Minimal [388]. Minutes [55, 143]. Mixtures [206]. Model [460, 220, 479, 219, 40, 204]. Modelling [459, 108]. Models [400, 285, 77, 248, 331, 306, 461, 45, 284, 88, 372, 368, 455, 388, 156, 473, 111]. Modern [156, 207, 510, 446, 272, 290]. Molecular [83]. Monitoring [176, 511]. Montgomery [88]. Monthly [120]. Moore [46]. Morbidity [165, 168]. Morphometrics [212]. Morrell [276, 505]. Morrison [156, 88, 134]. Morton [332]. Moser [153]. Most [188]. Mothballs [395]. Moving [301, 317]. Multi [203, 387]. Multi-Dimensional [387]. Multi-Variate [203]. Multidimensional [428, 315]. Multiple [201, 175]. Multivariate [255, 10, 84, 383, 252, 393, 269, 288, 476, 204, 302, 358, 212, 467, 420]. Municipal [86].

Must [353, 429]. Mutual [448]. Myers [445].

N [155, 22, 75, 450, 116, 108, 269]. N.I.E.S.R. [306]. Nachum [330]. Nain [390]. Narsingh [449]. National [541, 166, 481, 542]. Nationalized [20]. Nature [163]. Neave [472]. Nerlove [152]. Newell [90]. Newer [485]. Newman [126]. Newmark [446]. Newson [95]. Newton [438]. Neyman [123]. Nicholas [212]. Nievergeld [449]. Noise [104]. Non [416, 286, 37, 455, 177, 375, 273, 79]. Non-Lattice [286]. Non-Linear [455, 375]. Non-Numerical [273]. Non-Parametric [37, 177, 79]. Non-Sampling [416]. Nonparametric [211, 211, 463]. Normal [58]. Norton [59]. Nosal [444]. Note [98, 54]. Notice [536]. November [47, 125]. Number [70]. Numbers [85, 441, 516, 266]. Numeracy [495]. Numerical [273]. Nuptiality [356].

O [378, 242]. Obituary [226, 487, 500]. Objective [324]. Observations [497, 415, 305]. Obtaining [170]. O’Connor [392]. Odd [266]. Odell [126]. Office [149]. Official [535]. Oldham [91]. Olive [364]. One [370, 493, 203, 383, 387]. Open [465, 281]. Opening [404]. Operating [21]. Operations [29]. Opinion [121]. Optimal [128]. Optimisation [93, 329]. Order [58, 302]. Organizer [184]. Orientation [418]. Oswald [226]. Other [168]. Outlets [140]. Output [392]. Ownership [229].

P [474, 135, 23, 91, 77, 136, 46, 450, 126, 306, 441, 363, 10, 475, 59, 359, 361, 463, 464, 393, 329, 304, 467, 25]. Painless [441]. Paper [389, 496, 257]. Papers [306, 275, 192]. Parabolic [478]. Parametric [37, 177, 79]. Parsons [506]. Parthasarathy [526]. Partially [67, 293]. Particular [385]. Pass [96, 47, 125, 180]. Passes [369]. Patient [166]. Patil [135, 23, 136, 135, 77, 136]. Patterns [27]. Paul [130]. Pearson [114, 242, 475, 432, 123, 314]. RSS [520]. Tobin [219]. Penetration [256]. Percentage [499]. Percentages [441]. Performance [264]. Periods [341]. Personal [415]. Peter [488, 111, 159]. Phenomena [268]. Physical [136, 393]. Pitt [521]. Place [222]. Plackett [470]. Planning [36, 522, 156, 52]. Plotting [297]. Point [499]. Poisson [95]. Polly [116]. Polytechnic [354]. POPLAB [439]. Population [270, 168, 51, 52, 53, 250, 241]. Populations [481, 542, 221, 541, 170]. Portfolio [218]. Position [405]. Postal [415]. Poston [507]. Pour [60]. Power [497, 431, 347, 313, 205]. Practical [187, 341, 220, 433, 373, 108, 338, 209, 211]. Practice [166, 427, 376, 449, 276, 475, 13, 505]. Practised [353]. Preach [353]. Presentation [27]. Presenting [493]. President [82]. Press [448, 465]. Price [263]. Primer [361]. Principal [85, 86]. Principles [328, 522, 24, 46]. Probabilities [58]. Probability [74, 526, 311, 42, 297, 247, 129, 409, 444, 470, 303, 446, 9, 324, 421, 42, 114, 525, 330, 522, 450, 445, 523, 443, 56]. Probit [150]. Problem [513, 130, 527, 19, 238, 221]. Problems [198, 187, 186, 190, 95, 147, 289, 247, 363, 333, 107, 170, 95, 59, 471, 537, 375]. Procedures [407, 25]. Proceeding [213]. Proceedings [29]. Processes [301, 317, 410, 496, 56, 74, 473, 331]. Product [139]. Production [483]. Productivity [6]. Professor [487, 200]. Program [439]. Programming [151, 412, 375, 502]. Programs [237]. Project [340]. Projections [69]. Projects [480, 338]. Proof [525]. Properties [410]. Proportion [401, 357]. Proportions [327]. Prospects [190]. Provision [535]. Puri [79]. Purpose [163, 146].

Quality [468, 530, 177, 132]. Quantitative [238, 233, 59, 151, 519, 524]. Quasar [188]. Question [493]. Questionnaires [415]. Queueing [90]. Quick [144].

R [213, 360, 526, 43, 27, 326, 440, 506, 95, 306, 472, 72, 271, 276, 129, 212, 156, 92, 59, 76, 158, 274, 470, 466, 505, 151, 322, 361, 463, 464, 275, 113, 277, 304, 467, 392, 443, 134, 537, 25, 56]. R. [510]. Rabinovitch [330]. Radhakrishna [333]. Radical [353]. Raghavarao [107]. Rahman [22]. Ralston [7]. Ramachandran [333]. Random [135, 416, 74, 136, 206, 70, 85, 126, 56, 135, 136, 135, 77, 136]. Rank [73, 58]. Rao [333, 322]. Rate [494, 456, 458, 429]. Rates [176, 356, 327, 429]. Rather [529]. Rathie [450]. Ratio [208]. Rational [27]. Ratios [441]. Raymond [445, 523]. Readings [152, 59]. Real [359]. Recapture [368]. Reconciliation [141]. Record [169]. Reference [425, 385, 273]. Reflections [195, 192]. Regenerative [268]. Regime [106]. Regional [306]. Regions [78]. Regression [196, 298, 299, 111, 257, 374, 364, 305, 112]. Reingold [449]. Related [111]. Relevant [359]. Remarks [163, 456]. Reply [200]. Report [14, 185]. Reports [4, 31]. Representing [255]. Reprise [542]. Research [146, 29, 447, 492, 247, 158, 294, 6, 210, 117, 274, 109, 131, 420]. Residential [256]. Response [203, 383, 387]. Responses [428]. Responsibility [18]. Results [85]. Retail [140]. Revealing [499]. Review [439, 320, 321, 358, 376, 377, 379, 391, 474, 213, 26, 135, 23, 90, 360, 89, 231, 324, 325, 330, 74, 155, 78, 24, 291, 73, 91, 473, 526, 468, 521, 522, 22, 43, 27, 28, 29, 75, 77, 79, 136, 229, 326, 440, 412, 465, 328, 502, 331, 210, 130, 46, 450, 506, 353, 5, 11, 421, 520, 209, 94, 95, 211, 42, 126, 127, 157, 237, 58, 306, 115, 441, 445, 447, 449, 472, 527, 72, 116, 232, 271, 272, 276, 290, 153, 235, 305, 323, 363, 448, 503, 7, 10]. Review [45, 378, 234, 270, 241, 475, 12, 13, 57, 128, 129, 133, 148, 364, 486, 488, 292, 212, 240, 156, 238, 319, 87, 44, 132, 523, 60, 117, 92, 111, 233, 469, 332, 444, 471, 59, 76, 110, 114, 124, 149, 158, 274, 359, 470, 362, 108, 466, 504, 505, 501, 88, 294, 333, 327, 151, 322, 361, 463, 464, 393, 112, 275, 113, 152, 239, 329, 538, 109, 442, 485, 507, 508, 154, 159, 242, 277, 293, 380, 304, 107, 467, 268, 519, 150, 269, 93, 392, 131, 476, 443, 303, 236, 446, 525, 134, 175, 537]. Review [420, 524, 8, 6, 9, 25, 56, 375, 273]. Reviews [520]. Reyment [212]. Richard [157, 441, 525]. Ridge [196, 299, 257]. Risk [503, 533]. Rivals [510]. Robert [326, 294]. Robin [212]. Robust [70, 304]. Rodgers [238]. Roger [12, 148]. Rogers [304, 127]. Rohlf [326]. Role [342, 492, 234, 102, 168, 480, 53]. Ronald [445]. Roussas [303]. Roy [58]. Run [300]. Runyon [59, 441, 463, 464]. Russell [527].

S [321, 23, 522, 75, 7, 45, 13, 129, 114, 108, 112, 239, 242, 269, 375]. Safety [105]. Sales [67, 65]. Salve [395]. Sample [176, 58, 203, 383, 429, 386]. Sampling [416, 298, 531, 532, 223, 468]. Saturated [21]. Scale [120]. Schauer [359]. Schmalensee [360]. School [345]. Schools [399]. Science [136, 486, 109]. Sciences [135, 102, 391, 135, 472]. Scientific [135, 136, 231, 325, 210]. Scientists [445, 501, 56]. Scope [465]. Scope/Univariate [465]. Scottish [167, 86]. Seal [503]. Seasonal [310, 456, 458]. Seasonally [459]. Second [446]. Sections [256]. Segmentation [385]. Selected [192, 152]. Selection [374]. Sellers [440]. Semi [478]. Semi-Parabolic [478]. Sensitivity [347, 313]. Sequential [21, 142, 320]. Serial [410]. Series [460, 376, 379, 400, 389, 417, 474, 430, 459, 461, 378, 191, 455, 511, 249, 235, 323, 377].

Service [167, 367, 346, 541, 166, 481, 542, 40]. Setting [465, 67]. Seventy [366]. Seventy-Six [366]. Sewell [487, 500]. Sex [418]. Sexual [204]. Shackles [533]. Shapiro [45]. Share [220]. Sharpe [219]. Sharpe/Tobin [219]. Sheppard [234]. Shift [58]. Short [300]. Short-Run [300]. Showing [430]. Sibson [212]. Sides [493]. Signal [361]. Significance [134]. Silvey [321]. Simple [499]. Simulation [300, 157]. Since [292]. Sine [518]. Situation [203]. Six [366]. Sizes [176]. Skewed [221]. Slater [237]. Slee [130]. Slippage [144, 205]. Slope [511]. Small [41, 357]. Smith [271, 130]. Social [391, 135, 291, 331, 153, 486, 197, 256, 71, 59, 274]. Society [292, 380]. Sokal [326]. Solutions [537]. Solving [527, 130, 95]. Some [389, 496, 163, 497, 120, 218, 415, 147, 515, 494, 401, 516, 427, 38, 310, 20, 144, 205, 510, 195, 288, 541, 303, 257, 316]. Source [168]. Sources [520, 291]. Spaces [281]. Spatial [473, 287, 284, 249, 250, 316]. Special [425, 273]. Specific [356]. Speckman [154]. Spectral [454, 254, 189]. Spirits [178]. Sprent [111]. Squares [438]. SRS [357]. SSRC [520]. SSRC/RSS [520]. Standard [228, 310]. Standardizing [197]. States [332]. Statistic [513, 515, 432, 314]. Statistical [349, 391, 2, 41, 185, 36, 231, 330, 155, 286, 146, 187, 79, 407, 186, 529, 218, 46, 353, 520, 95, 211, 157, 289, 227, 235, 128, 217, 480, 531, 530, 76, 149, 38, 322, 355, 195, 339, 443, 532, 223, 249, 204, 8, 9, 25, 45, 321, 78, 521, 506, 271, 448, 133, 158, 501, 327, 109, 159, 6]. Statistician [349, 370, 174, 232, 102, 532]. Statisticians [213, 492, 341, 242, 82, 96, 55, 47, 143, 125, 180]. Statistics [379, 337, 535, 28, 412, 465, 95, 343, 441, 445, 340, 272, 397, 503, 45, 129, 148, 364, 486, 399, 44, 59, 114, 470, 188, 466, 413, 173, 345, 282, 20, 463, 464, 367, 269, 338, 435, 222, 245, 103, 414, 122, 250, 302, 95, 59, 320, 291, 22, 43, 27, 440, 450, 209, 211, 72, 116, 488, 240, 471, 333, 303, 537, 26, 5, 447, 472, 276, 12, 57, 469, 359, 504, 505, 277, 380, 446]. Statute [407]. Step [511]. Stewart [507]. Stochastic [331, 88]. Strategies [425]. Stratified [479]. Structure [83, 42]. Structures [77]. Stuart [379, 28]. Students [354, 95, 521, 486]. Studies [44, 114, 470]. Study [124, 205, 386, 103]. Subjective [67]. Subroutine [208]. Sum [13]. Summing [170]. Sums [206]. Survey [27, 120, 385, 121, 304, 153]. Surveys [168]. Survival [503]. Survivor [37]. Symbols [352]. Symposium [185]. System [39, 67, 530, 479]. System-Theoretic [530]. Systems [169, 162, 163, 167, 228, 310].

T [43, 126, 72, 235, 108, 113, 380]. Table [387, 361]. Tables [391, 436, 472, 373, 242, 428, 315, 293]. Tabulated [538]. Taking [54]. Tang [522]. Tanis [443]. Tanks [130]. Targets [67]. Taxanomy [212]. Teaching [337, 353, 354, 399, 345, 339, 367, 338]. Technique [256, 295, 265]. Techniques [79, 229, 255, 297, 189, 68, 476, 142, 6, 273, 93]. Technological [6]. Technology [504]. Technques [46]. Telephone [256]. Temperature [106]. Temporal [287]. Term [69]. Terms [76]. Test [515, 433, 514, 134]. Testing [139, 123]. Tests [13, 373, 144, 205]. Thad [466]. Thampson [238]. Theil [328]. Their [535, 146, 306]. Theoretic [198, 530]. Theoretical [537]. Theortical [22]. Theory [376, 379, 324, 28, 421, 42, 449, 276, 13, 427, 505, 507, 9, 123, 90, 94, 503, 361]. Think [130]. Thinking [529]. Third [213, 103]. Thomas [74]. Thomasian [42]. Thoughts [416]. Three [369]. Tim [507]. Time [460, 376, 379, 410, 400, 389, 417, 430, 459, 461, 235, 191, 455, 511, 377, 474, 323, 378]. Time- [379]. Time-Series [430, 377].

Tolerance [78]. Tomassone [213]. Tool [262]. Tools [227]. Top [369]. Topics [270, 111]. Toy [262]. Trades [370]. Training [492, 341]. Transformations [497, 431, 312]. Transients [511]. Transition [372]. Trees [491]. Trend [69, 264]. Trends [71, 292]. Trial [515, 437]. Trials [512, 289, 433]. Truth [231]. Tufte [59]. Tukey [304]. Twenty [14]. Twenty-First [14]. Two [262, 70, 58, 68, 437, 293, 142]. Two-Armed [437]. Two-Associate-Class [293]. Two-Edged [262]. Two-Sample [58].

U.K [234]. U.S. [456, 458, 53]. Unbiased [357]. Uncertainty [258, 318, 11]. Understanding [146, 156]. Undertaking [20]. Unemployment [306, 494, 456, 458]. Unequal [298]. Unfamiliar [513]. Unified [167]. United [520, 332, 124]. Univariate [75, 465, 204]. Universities [367, 338]. University [465, 355]. Upton [538]. Urban [106, 447]. Use [120, 115, 256, 282, 388]. Used [229]. User [173]. Uses [297, 20, 485]. Using [176]. Utilisation [165].

V [89, 471, 333, 443, 537]. Vajda [375]. Vale [395]. Validity [40]. Value [535]. Variable [374]. Variables [188]. Variance [364, 347, 313]. Variate [203]. Variates [126]. Variation [430, 284]. Varying [176]. Venkatesan [210]. Vice [82]. Vice-President [82]. Vie [60]. View [535, 196, 167, 299, 413]. Viewpoint [283, 531, 532]. Virginia [364]. Visual [27, 255]. Vol [379, 135, 522, 136, 520]. Volk [45]. Volume [99, 119, 181, 224, 259, 334, 381, 423, 490, 539, 28, 270, 470, 275, 242]. Voting [86].

W [391, 23, 326, 421, 520, 209, 94, 95, 211, 72, 272, 235, 45, 59, 110, 362, 152, 154, 159, 380, 304, 392, 6, 25]. Walpole [445]. Wastage [38, 372]. Watch [429]. Watson [108]. Watson-Gandy [108]. Weapon [262]. Weather [258, 318]. Weaver [421]. Weeden [306]. Weighted [438]. Wetherill [320, 468]. Whittle [329]. Wichern [496]. Wide [281]. Wilf [7]. Wilfrid [8]. Willard [293]. William [42, 238, 76, 88]. Williams [95, 232, 271, 290]. Williamson [389, 241]. Wilson [522]. Winning [441]. Within [542, 481, 541]. Witness [406, 405]. Wolock [154]. Women [264]. Wonnacott [43, 113, 43, 113]. Worcester [294]. Word [309]. Wording [493]. Work [340, 135, 136]. Workers [158]. Written [199].

X [459]. X-11 [459].

Yannoulis [391]. Yates [192]. Year [14, 354]. Yearbook [448]. Yu [471, 333].

Z [391].

References

Bissell:1970:ADB

[1] A. F. Bissell. Analysis of data based on incident counts. Journal of the Royal Statistical Society. Series D (The Statistician), 19(3):215–247, ???? 1970. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                             www.jstor.org/stable/2986835.

Anonymous:1970:SA

[2] Anonymous. Statistical algorithms. Journal of the Royal Statistical Society. Series D (The Statistician), 19(3): 249–254, ???? 1970. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                             tronic). URL http://www.jstor.org/
                                             stable/2986836.

Nelder:1970:CDA

[3] J. A. Nelder, B. E. Cooper, J. M. Craddock, J. C. Gower, P. J. Harrison, and I. D. Hill. The construction and description of algorithms. Journal of the Royal Statistical Society. Series D (The Statistician), 19(3):255–260, ???? 1970. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986837.

Anonymous:1970:EERa

[4] Anonymous. 1969 examinations examiners’ reports. Journal of the Royal Statistical Society. Series D (The Statistician), 19(3):261–345, ???? 1970. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986838.

Elton:1970:BRB

[5] H. D. Elton. Book review: Statistics in the Computer Age, by J. M. Craddock. Journal of the Royal Statistical Society. Series D (The Statistician), 19(3):347, ???? 1970. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986839.

Westlake:1970:BRBa

[6] A. J. Westlake. Book review: Statistical Techniques in Technological Research — An Aid to Research Productivity, by W. E. Duckworth. Journal of the Royal Statistical Society. Series D (The Statistician), 19(3):348–350, ???? 1970. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986840.

Hawkins:1970:BRBa

[7] Douglas M. Hawkins. Book review: Mathematical Methods for Digital Computers, by A. Ralston; H. S. Wilf. Journal of the Royal Statistical Society. Series D (The Statistician), 19 (3):350, ???? 1970. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986841.

Ward:1970:BRB

[8] D. H. Ward. Book review: Introduction to Statistical Analysis, by Wilfrid J. Dixon; Frank J. Massey. Journal of the Royal Statistical Society. Series D (The Statistician), 19(3):351–352, ???? 1970. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986842.

Westlake:1970:BRBb

[9] A. J. Westlake. Book review: Introduction to Probability Theory and Statistical Inference, by Harold J. Larson. Journal of the Royal Statistical Society. Series D (The Statistician), 19(3): 352–353, ???? 1970. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986843.

Hawkins:1970:BRBb

[10] Douglas M. Hawkins. Book review: Elements of Continuous Multivariate Analysis, by A. P. Dempster. Journal of the Royal Statistical Society. Series D (The Statistician), 19(3):353–354, ???? 1970. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986844.

Fawthrop:1970:BRB

[11] J. Fawthrop. Book review: Uncertainty and Estimation in Economics, by D. G. Champernowne. Journal of the Royal Statistical Society. Series D (The Statistician), 19(3):354–356, ???? 1970. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986845.

Huitson:1970:BRBa

[12] Alan Huitson. Book review: Statistics: A Computer Approach, by Roger L. Burford. Journal of the Royal Statistical Society. Series D (The Statistician), 19 (3):356–357, ???? 1970. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986846.

Huitson:1970:BRBb

[13] A. Huitson. Book review: Cumulative Sum Tests: Theory and Practice, by C. S. van Dobben de Bruyn. Journal of the Royal Statistical Society. Series D (The Statistician), 19(3):357, ???? 1970. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986847.

Anonymous:1970:RCY

[14] Anonymous. Report of Council for the year 1969 to the Twenty-First Annual General Meeting. Journal of the Royal Statistical Society. Series D (The Statistician), 19(3):358–367, ???? 1970. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986848.

Anonymous:1970:Ea

[15] Anonymous. Education. Journal of the Royal Statistical Society. Series D (The Statistician), 19(3):368–374, ???? 1970. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2986849.

Anonymous:1970:BMa

[16] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 19(3):??, ???? 1970. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2986850.

Anonymous:1970:FMa

[17] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 19(3):??, ???? 1970. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2986834.

Anonymous:1970:ERI

[18] Anonymous. Editorial: The responsibility of the Institute. Journal of the Royal Statistical Society. Series D (The Statistician), 19(4):375–376, ???? 1970. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987370.

Gupta:1970:ADA

[19] S. K. Sen Gupta. Application of discriminatory analysis to an engineering problem. Journal of the Royal Statistical Society. Series D (The Statistician), 19(4):377–389, ???? 1970. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987371.

Nabb:1970:SUS

[20] H. Nabb. Some uses of statistics in a nationalized gas undertaking. Journal of the Royal Statistical Society. Series D (The Statistician), 19(4):391–402, ???? 1970. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987372.

Shahani:1970:SES

[21] A. K. Shahani. A saturated experiment in sequential determination of operating conditions. Journal of the Royal Statistical Society. Series D (The Statistician), 19(4):403–408, ???? 1970. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987373.

Blackmore:1970:BRB

[22] R. W. Blackmore. Book review: A Course in Theortical Statistics, by N. A. Rahman. Journal of the Royal Statistical Society. Series D (The Statistician), 19(4):409, ???? 1970. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987374.

Armitage:1970:BRB

[23] P. Armitage. Book review: A Dictionary and Bibliography of Discrete Distributions, by G. P. Patil; S. W. Joshi. Journal of the Royal Statistical Society. Series D (The Statistician), 19(4): 409–410, ???? 1970. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987375.

Benjamin:1970:BRB

[24] B. Benjamin. Book review: Principles of Demography, by Donald J. Bogue. Journal of the Royal Statistical Society. Series D (The Statistician), 19(4):

410–411, ???? 1970. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987376.

Westlake:1970:BRBc

[25] A. J. Westlake. Book review: Introduction to Statistical Procedures: With Computer Exercises, by P. R. Lohnes; W. W. Cooley. Journal of the Royal Statistical Society. Series D (The Statistician), 19(4):411–413, ???? 1970. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987377.

Anonymous:1970:BRB

[26] Anonymous. Book review: Statistics for Mathematicians — An Introduction, by D. J. Finney. Journal of the Royal Statistical Society. Series D (The Statistician), 19(4):413–414, ???? 1970. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987378.

Buckland:1970:BRBa

[27] W. R. Buckland. Book review: Graphical Rational Patterns: A New Approach to Graphical Presentation of Statistics, by R. Bachi; Diagrams: A Visual Survey of Graphs, Maps, Charts and Diagrams, by A. Lockwood. Journal of the Royal Statistical Society. Series D (The Statistician), 19(4):414–415, ???? 1970. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987379.

Buckland:1970:BRBb

[28] W. R. Buckland. Book review: Advanced Theory of Statistics Volume 1, by M. G. Kendall; A. Stuart. Journal of the Royal Statistical Society. Series D (The Statistician), 19(4):415–416, ???? 1970. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987380.

Buckland:1970:BRBc

[29] W. R. Buckland. Book review: 4th International Conference in Operations Research, Proceedings, by D. B. Hertz; J. Melese. Journal of the Royal Statistical Society. Series D (The Statistician), 19(4):417–418, ???? 1970. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987381.

Anonymous:1970:Eb

[30] Anonymous. Education. Journal of the Royal Statistical Society. Series D (The Statistician), 19(4):419–426, ???? 1970. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987382.

Anonymous:1970:EERb

[31] Anonymous. 1970 examinations examiners reports. Journal of the Royal Statistical Society. Series D (The Statistician), 19(4):427–503, ???? 1970. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987383.

Anonymous:1970:BMb

[32] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 19(4):504, ???? 1970. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987384.

Anonymous:1970:FMb

[33] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 19(4):??, ???? 1970. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987369.

Anonymous:1971:FMa

[34] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 20(1):1, March 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987001.

Anonymous:1971:Ea

[35] Anonymous. Editorial. Journal of the Royal Statistical Society. Series D (The Statistician), 20(1):2, March 1971. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987002.

Bartholomew:1971:SAM

[36] D. J. Bartholomew. The statistical approach to manpower planning. Journal of the Royal Statistical Society. Series D (The Statistician), 20(1):3–26, March 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2987003.

Forbes:1971:NPM

[37] A. F. Forbes. Non-parametric methods of estimating the survivor function. Journal of the Royal Statistical Society. Series D (The Statistician), 20 (1):27–52, March 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987004.

Marshall:1971:SSM

[38] M. L. Marshall. Some statistical methods for forecasting wastage. Journal of the Royal Statistical Society. Series D (The Statistician), 20(1):53–68, March 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987005.

Butlers:1971:AFM

[39] A. D. Butlers. An analysis of flows in a manpower system. Journal of the Royal Statistical Society. Series D (The Statistician), 20(1):69–84, March 1971. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987006.

Sales:1971:VMC

[40] Pauline Sales. The validity of the Markov chain model for a class of the Civil Service. Journal of the Royal Statistical Society. Series D (The Statistician), 20(1):85–110, March 1971. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987007.

Anonymous:1971:SES

[41] Anonymous. Statistical enquiries and the small business. Journal of the Royal Statistical Society. Series D (The Statistician), 20(1):111–115, March 1971. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987008.

Gani:1971:BRB

[42] J. Gani. Book review: Concepts of Probability, by William C. Guenther; The Structure of Probability Theory with Applications, by Aram J. Thomasian; Probability. by Leo Breiman. Journal of the Royal Statistical Society. Series D (The Statistician), 20(1):116–117, March 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987009.

Blackmore:1971:BRB

[43] R. W. Blackmore. Book review: Introductory Statistics, by T. H. Wonnacott; R. J. Wonnacott. Journal of the Royal Statistical Society. Series D (The Statistician), 20(1):117, March 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987010.

Kneller:1971:BRB

[44] Pamela G. Kneller. Book review: Applied Statistics for Management Studies, by David Croft. Journal of the Royal Statistical Society. Series D (The Statistician), 20(1):118, March 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987011.

Hawkins:1971:BRB

[45] D. Hawkins. Book review: Applied Statistics for Engineers, by W. Volk; Statistical Models in Engineering, by G. J. Hahn; S. S. Shapiro. Journal of the Royal Statistical Society. Series D (The Statistician), 20(1):118–119, March 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987012.

Duff:1971:BRB

[46] R. J. Duff. Book review: Principles of Statistical Technques, by P. G. Moore.

Journal of the Royal Statistical Society. Series D (The Statistician), 20(1): 119–120, March 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987013.

Anonymous:1971:ISN

[47] Anonymous. The Institute of Statisticians November 1970 examinations pass list. Journal of the Royal Statistical Society. Series D (The Statistician), 20 (1):120, March 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987014.

Anonymous:1971:BMa

[48] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 20(1):??, March 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987015.

Anonymous:1971:FMb

[49] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 20(2):1, June 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987030.

Anonymous:1971:Eb

[50] Anonymous. Editorial. Journal of the Royal Statistical Society. Series D (The Statistician), 20(2):2, June 1971. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987031.

Newman:1971:CPB

[51] Dennis Newman. The census of population in Britain. Journal of the Royal Statistical Society. Series D (The Statistician), 20(2):3–14, June 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2987032.

Newman:1971:PCP

[52] Dennis Newman. The planning of the 1971 census of population. Journal of the Royal Statistical Society. Series D (The Statistician), 20(2):15–31, June 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2987033.

Waksberg:1971:REU

[53] J. Waksberg and W. M. Perkins. The role of evaluation in U.S. censuses of population and housing. Journal of the Royal Statistical Society. Series D (The Statistician), 20(2):33–46, June 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987034.

Lury:1971:NCT

[54] D. A. Lury. A note on census-taking during 1955–64. Journal of the Royal Statistical Society. Series D (The Statistician), 20(2):47–52, June 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2987035.

Anonymous:1971:ISM

[55] Anonymous. The Institute of Statisticians minutes of the 21st Annual General Meeting. Journal of the Royal Statistical Society. Series D (The Statistician), 20(2):53–55, June 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987036.

Westlake:1971:BRB

[56] A. J. Westlake. Book review: Probability and Random Processes for Engineers and Scientists, by A. B. Clarke; R. L. Disney. Journal of the Royal Statistical Society. Series D (The Statistician), 20(2):57, June 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987037.

Huitson:1971:BRB

[57] A. Huitson. Book review: Statistics for Experimentalists, by B. E. Cooper. Journal of the Royal Statistical Society. Series D (The Statistician), 20(2):58, June 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987038.

Gent:1971:BRB

[58] C. R. Gent. Book review: Rank Order Probabilities: Two-Sample Normal Shift Alternatives, by Roy C. Milton. Journal of the Royal Statistical Society. Series D (The Statistician), 20 (2):58–59, June 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987039.

Lury:1971:BRBa

[59] D. A. Lury. Book review: Elementary Statistics for Economics and Business, by E. Mansfield; W. W. Norton; The Quantitative Analysis of Social Problems, by E. R. Tufte; Readings in Statistics, by A. Haber; R. P. Runyon; P. Badia. Journal of the Royal Statistical Society. Series D (The Statistician), 20(2):59–60, June 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987040.

L:1971:BRB

[60] D. A. L. Book review: Documents Pour l’Élaboration d’Indices due Coût de la Vie, by J. Fourastie. Journal of the Royal Statistical Society. Series D (The Statistician), 20(2):60, June 1971. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987041.

Anonymous:1971:BMb

[61] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 20(2):??, June 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987042.

Anonymous:1971:FMc

[62] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 20(3):1, September 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2986790.

Anonymous:1971:Ec

[63] Anonymous. Editorial. Journal of the Royal Statistical Society. Series D (The Statistician), 20(3):2, September 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2986791.

Rodger:1971:FAG

[64] A. A. S. Rodger, W. J. Rawlings, and J. M. Simpson. Forecasting in an area gas board. Journal of the Royal Statistical Society. Series D (The Statistician), 20(3):3–14, September 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2986792.

Myers:1971:FES

[65] C. L. Myers. Forecasting electricity sales. Journal of the Royal Statistical Society. Series D (The Statistician), 20(3): 15–22, September 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986793.

Gearing:1971:FD

[66] H. W. G. Gearing. Forecasting demand. Journal of the Royal Statistical Society. Series D (The Statistician), 20(3): 23–28, September 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986794.

Coates:1971:PSS

[67] R. G. Coates. A partially subjective system for setting sales targets. Journal of the Royal Statistical Society. Series D (The Statistician), 20(3): 29–34, September 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986795.

Marchant:1971:CTF

[68] L. J. Marchant and D. J. Hockley. A comparison of two forecasting techniques. Journal of the Royal Statistical Society. Series D (The Statistician), 20(3):35–44, September 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2986796.

Dugdale:1971:LTT

[69] Madge Dugdale. Long term trend projections as an aid to management. Journal of the Royal Statistical Society. Series D (The Statistician), 20(3): 45–54, September 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986797.

Craddock:1971:TRM

[70] J. M. Craddock and Sheila A. Farmer. Two robust methods of random number generation. Journal of the Royal Statistical Society. Series D (The Statistician), 20(3):55–66, September 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986798.

Ilersic:1971:STC

[71] A. R. Ilersic and R. A. Pluck. Social trends: A comment. Journal of the Royal Statistical Society. Series D (The Statistician), 20(3):67–72, September 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2986799.

Grimshaw:1971:BRB

[72] J. J. Grimshaw. Book review: Managing with Statistics, by T. W. Lewis; R. A. Fox. Journal of the Royal Statistical Society. Series D (The Statistician), 20(3):73, September 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986800.

Bevan:1971:BRBa

[73] J. M. Bevan. Book review: Rank Correlation Methods, by M. G. Kendall. Journal of the Royal Statistical Society. Series D (The Statistician), 20(3): 74, September 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986801.

Basilevsky:1971:BRB

[74] A. Basilevsky. Book review: An Introduction to Applied Probability and Random Processes, by J. B. Thomas. Journal of the Royal Statistical Society. Series D (The Statistician), 20(3): 74–75, September 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986802.

Buckland:1971:BRBa

[75] W. R. Buckland. Book review: Discrete Distributions; Continuous Univariate Distributions I and II, by N. L. Johnson; S. Kotz. Journal of the Royal Statistical Society. Series D (The Statistician), 20(3):75, September 1971. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986803.

Lury:1971:BRBb

[76] D. A. Lury. Book review: A Dictionary of Statistical Terms, by Maurice G. Kendall; William R. Buckland. Journal of the Royal Statistical Society. Series D (The Statistician), 20(3): 76, September 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986804.

Buckland:1971:BRBb

[77] W. R. Buckland. Book review: Random Counts in Models and Structures, by G. P. Patil. Journal of the Royal Statistical Society. Series D (The Statistician), 20(3):76–77, September 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2986805.

Bassett:1971:BRB

[78] Eryl Bassett. Book review: Statistical Tolerance Regions: Classical and Bayesian, by I. Guttman. Journal of the Royal Statistical Society. Series D (The Statistician), 20(3):77–78, September 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2986806.

Buckland:1971:BRBc

[79] W. R. Buckland. Book review: Non-Parametric Techniques in Statistical Inference, by M. L. Puri. Journal of the Royal Statistical Society. Series D (The Statistician), 20(3):78, September 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2986807.

Anonymous:1971:BMc

[80] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 20(3):??, September 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2986808.

Anonymous:1971:FMd

[81] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 20(4):1, December 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2986982.

Anonymous:1971:ISA

[82] Anonymous. The Institute of Statisticians: Appointment of new President and Vice-President. Journal of the Royal Statistical Society. Series D (The Statistician), 20(4):2, December 1971. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986983.

Brown:1971:KCB

[83] P. J. Brown. Keys to correlate biological activity with molecular structure of chemical compounds. Journal of the Royal Statistical Society. Series D (The Statistician), 20(4):3– 49, December 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986984.

Krzanowski:1971:ABC

[84] W. J. Krzanowski. The algebraic basis of classical multivariate methods. Journal of the Royal Statistical Society. Series D (The Statistician), 20(4): 51–61, December 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986985.

Farmer:1971:IRP

[85] S. A. Farmer. An investigation into the results of principal component analysis of data derived from random numbers. Journal of the Royal Statistical Society. Series D (The Statistician), 20(4):

63–72, December 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2986986.

Lincoln:1971:APC

[86] S. V. Lincoln, Anthony Piepe, and R. Prior. An application of principal components analysis to voting in Scottish municipal elections 1967–9. Journal of the Royal Statistical Society. Series D (The Statistician), 20(4):73– 88, December 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2986987.

Kirk:1971:BRB

[87] J. H. Kirk. Book review: Experimental Marketing, by E. J. Davis. Journal of the Royal Statistical Society. Series D (The Statistician), 20(4):89, December 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2986988.

Marchant:1971:BRB

[88] L. Marchant. Book review: Stochastic Models of Buying Behaviour, by William F. Massy; David B. Montgomery; David G. Morrison. Journal of the Royal Statistical Society. Series D (The Statistician), 20(4):90, December 1971. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2986989.

Bartholomew:1971:BRB

[89] D. J. Bartholomew. Book review: Making Decisions, by D. V. Lindley. Journal of the Royal Statistical Society. Series D (The Statistician), 20(4): 90–92, December 1971. CODEN ????

ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986990.

Ashton:1971:BRB

[90] W. D. Ashton. Book review: Applications of Queueing Theory, by G. F. Newell. Journal of the Royal Statistical Society. Series D (The Statistician), 20(4):92–93, December 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2986991.

Bevan:1971:BRBb

[91] J. M. Bevan. Book review: Measurement in Medicine, by P. D. Oldham. Journal of the Royal Statistical Society. Series D (The Statistician), 20(4): 93–94, December 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986992.

Lancaster:1971:BRB

[92] A. Lancaster. Book review: Analysis of Binary Data, by D. R. Cox. Journal of the Royal Statistical Society. Series D (The Statistician), 20 (4):94, December 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986993.

Saalmans:1971:BRB

[93] Peter D. Saalmans. Book review: Introduction to Optimisation Techniques, by Masanao Aoki. Journal of the Royal Statistical Society. Series D (The Statistician), 20(4):95, December 1971. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986994.

Fuller:1971:BRBa

[94] M. F. Fuller. Book review: Elements of Decision Theory, by B. W. Lindgren. Journal of the Royal Statistical Society. Series D (The Statistician), 20(4): 96–97, December 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986995.

Fuller:1971:BRBb

[95] M. F. Fuller. Book review: Solving Problems in Statistics II, by J. Aitchison; The Statistical Argument, by R. M. Cormack; Advanced General Statistics, by B. C. Erricker; Statistics — Definitions and Formulae for Students, by R. W. Hiorns; The Language of Basic Statistics, by John Newson; Michael Matthews; Problems in Statistics — The Poisson and Exponential Distributions, by K. Williams. Journal of the Royal Statistical Society. Series D (The Statistician), 20(4):97–98, December 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986996.

Anonymous:1971:ISJ

[96] Anonymous. The Institute of Statisticians June 1971 examinations pass list. Journal of the Royal Statistical Society. Series D (The Statistician), 20(4): 99–103, December 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986997.

Anonymous:1971:A

[97] Anonymous. [announcements]. Journal of the Royal Statistical Society. Series D (The Statistician), 20(4):104– 105, December 1971. CODEN ????

ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986998.

Anonymous:1971:EN

[98] Anonymous. Editor’s note. Journal of the Royal Statistical Society. Series D (The Statistician), 20(4):106, December 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2986999.

Anonymous:1971:VI

[99] Anonymous. Volume information. Journal of the Royal Statistical Society. Series D (The Statistician), 20(4):107– 108, December 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986981.

Anonymous:1971:BMd

[100] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 20(4):??, December 1971. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987000.

Anonymous:1972:FMa

[101] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 21(1):1–2, March 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2986852.

Jeffers:1972:SRE

[102] J. N. R. Jeffers. The Statistician’s role in the environmental sciences. Journal of the Royal Statistical Society. Series D (The Statistician), 21(1):3–17, March 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2986853.

Thompson:1972:SET

[103] F. P. Thompson. Statistics and the environment — the Third London Airport Study. Journal of the Royal Statistical Society. Series D (The Statistician), 21 (1):19–30, March 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986854.

Flowerdew:1972:CAN

[104] A. D. J. Flowerdew. The cost of airport noise. Journal of the Royal Statistical Society. Series D (The Statistician), 21 (1):31–46, March 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986855.

Kirkby:1972:ASF

[105] W. W. Kirkby, G. W. A. Salmond, and K. W. Davies. The assessment for safety of food additives. Journal of the Royal Statistical Society. Series D (The Statistician), 21(1):47–61, March 1972. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986856.

Craddock:1972:UDE

[106] J. M. Craddock. Urban development and its effects on the local temperature regime. Journal of the Royal Statistical Society. Series D (The Statistician), 21(1):63–75, March 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986857.

Raghavarao:1972:CAE

[107] Damaraju Raghavarao and D. A.

Preece. Combinatorial analysis and experimental design: A review of Constructions and Combinatorial Problems in Design of Experiments, by Damaraju Raghavarao. Journal of the Royal Statistical Society. Series D (The Statistician), 21(1):77–87, March 1972. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986858.

Maddison:1972:BRB

[108] A. E. Maddison. Book review: Distribution Management: Mathematical Modelling and Practical Analysis, by S. Eilon; C. D. T. Watson-Gandy; N. Christofides. Journal of the Royal Statistical Society. Series D (The Statistician), 21(1):89–90, March 1972. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986859.

Pluck:1972:BRB

[109] R. A. Pluck. Book review: Statistical Methods for Behavioural Science Research, by L. A. Marascuilo. Journal of the Royal Statistical Society. Series D (The Statistician), 21(1):90, March 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2986860.

Lury:1972:BRBa

[110] D. A. Lury. Book review: Biological Aspects of Demography, by W. Brass. Journal of the Royal Statistical Society. Series D (The Statistician), 21 (1):90–91, March 1972. CODEN ????

ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986861.

Lancaster:1972:BRB

[111] A. Lancaster. Book review: Models in Regression and Related Topics, by Peter Sprent. Journal of the Royal Statistical Society. Series D (The Statistician), 21(1):91–92, March 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986862.

Pack:1972:BRB

[112] A. D. Pack. Book review: Regression and Econometric Methods, by David S. Huang. Journal of the Royal Statistical Society. Series D (The Statistician), 21(1):92–93, March 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986863.

Phillips:1972:BRBa

[113] Garry Phillips. Book review: Econometrics, by R. J. Wonnacott; T. H. Wonnacott. Journal of the Royal Statistical Society. Series D (The Statistician), 21 (1):93–94, March 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986864.

Lury:1972:BRBb

[114] D. A. Lury. Book review: Studies in the History of Statistics and Probability, by E. S. Pearson; M. G. Kendall. Journal of the Royal Statistical Society. Series D (The Statistician), 21 (1):95, March 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986865.

Goy:1972:BRB

[115] J. R. Goy. Book review: The Use of Economics Literature, by John Fletcher. Journal of the Royal Statistical Society. Series D (The Statistician), 21 (1):95–96, March 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986866.

Grimshaw:1972:BRB

[116] J. J. Grimshaw. Book review: A First Course in Statistics, by F. N. David; Bancroft’s Introduction to Biostatistics, by Johannes Ipsen; Polly Feigl. Journal of the Royal Statistical Society. Series D (The Statistician), 21 (1):96–97, March 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986867.

Lainton:1972:BRB

[117] K. Lainton. Book review: Leading Cases in Market Research, by Max K. Adler. Journal of the Royal Statistical Society. Series D (The Statistician), 21 (1):97–98, March 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986868.

Anonymous:1972:BMa

[118] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 21(1):??, March 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2986869.

Anonymous:1972:VI

[119] Anonymous. Volume information. Journal of the Royal Statistical Society. Series D (The Statistician), 21(1):??, March 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2986851.

Daly:1972:SEL

[120] Joseph F. Daly, Harvey Kailin, Ralph S. Woodruff, and Max Shor. Some experience in the large scale use of administrative lists in a monthly business survey. Journal of the Royal Statistical Society. Series D (The Statistician), 21 (2):99–113, June 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987321.

Harris:1972:OMS

[121] P. T. Harris. Opinion and market survey methodology. Journal of the Royal Statistical Society. Series D (The Statistician), 21(2):114–125, June 1972. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987322.

Williams:1972:SC

[122] Ken Williams. Statistics and communication. Journal of the Royal Statistical Society. Series D (The Statistician), 21 (2):127–128, June 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987323.

Williams:1972:CNP

[123] K. Williams. A criticism of the Neyman– Pearson theory of testing. Journal of the Royal Statistical Society. Series D (The Statistician), 21(2):128–131, June 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987324.

Lury:1972:BRBc

[124] D. A. Lury. Book review: The Econometric Study of the United Kingdom. Journal of the Royal Statistical Society. Series D (The Statistician), 21(2): 132–140, June 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987325.

Anonymous:1972:ISN

[125] Anonymous. Institute of Statisticians November 1971 examination pass list. Journal of the Royal Statistical Society. Series D (The Statistician), 21(2): 141–142, June 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987326.

Garside:1972:BRBa

[126] M. J. Garside. Book review: The Generation of Random Variates, by T. G. Newman; P. L. Odell. Journal of the Royal Statistical Society. Series D (The Statistician), 21(2):143, June 1972. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987327.

Garside:1972:BRBb

[127] M. J. Garside. Book review: Basic Fortran 4, by B. J. Dowding; C. B. Rogers. Journal of the Royal Statistical Society. Series D (The Statistician), 21(2):143–144, June 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987328.

Huitson:1972:BRBa

[128] A. Huitson. Book review: Optimal Statistical Decisions, by M. H. DeGroot. Journal of the Royal Statistical Society. Series D (The Statistician), 21(2):144, June 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987329.

Huitson:1972:BRBb

[129] A. Huitson. Book review: Handbook of Probability and Statistics. 2nd Ed, by R. S. Burington; D. C. May. Journal of the Royal Statistical Society. Series D (The Statistician), 21(2):144, June 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987330.

Dowie:1972:BRB

[130] J. A. Dowie. Book review: Think Tanks and Problem Solving, by Paul I. Slee Smith. Journal of the Royal Statistical Society. Series D (The Statistician), 21(2):145, June 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987331.

Seymour-Ure:1972:BRB

[131] Colin Seymour-Ure. Book review: Content Analysis in Communication Research, by Bernard Berelson. Journal of the Royal Statistical Society. Series D (The Statistician), 21(2):145, June 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987332.

Kneller:1972:BRB

[132] P. G. Kneller. Book review: Quality Control Bookwork, by David J.

Desmond. Journal of the Royal Statistical Society. Series D (The Statistician), 21(2):146, June 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987333.

Huitson:1972:BRBc

[133] A. Huitson. Book review: Statistical Analysis for Managerial Decisions, by J. C. G. Boot; E. B. Cox. Journal of the Royal Statistical Society. Series D (The Statistician), 21(2):147, June 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987334.

Taylor:1972:BRB

[134] P. S. Taylor. Book review: The Significance Test Controversy, by R. E. Henkel; D. E. Morrison. Journal of the Royal Statistical Society. Series D (The Statistician), 21(2):147–149, June 1972. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987335.

Anonymous:1972:BRB

[135] Anonymous. Book review: Random Counts in Biomedical and Social Sciences, by G. P. Patil; Random Counts in Biomedical and Social Sciences: Vol. I. Vol. II. Random Counts in Scientific Work, by G. P. Patil. Journal of the Royal Statistical Society. Series D (The Statistician), 21(2):149, June 1972. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987336.

Buckland:1972:BRB

[136] W. R. Buckland. Book review: Random Counts in Physical Science, Geo Science and Business, by G. P. Patil; Random Counts in Physical Science, Geo Science and Business: Vol III, Random Counts in Scientific Work, by G. P. Patil. Journal of the Royal Statistical Society. Series D (The Statistician), 21(2):149–150, June 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987337.

Anonymous:1972:BMb

[137] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 21(2):??, June 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987338.

Anonymous:1972:FMb

[138] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 21(2):??, June 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987320.

Boyd:1972:PT

[139] K. T. Boyd. Product testing. Journal of the Royal Statistical Society. Series D (The Statistician), 21(3):151– 159, September 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2986678.

Sparks:1972:MER

[140] D. N. Sparks and R. K. Landells. Marketing experiments in retail grocery outlets — the Contimart Marketing Laboratory. Journal of the Royal Statistical Society. Series D (The Statistician), 21 (3):161–167, September 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2986679.

Boggis:1972:DCR

[141] J. G. Boggis. Data conflict and reconciliation. Journal of the Royal Statistical Society. Series D (The Statistician), 21 (3):169–180, September 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2986680.

Thompson:1972:SDT

[142] V. R. Thompson. Sequential dichotomisation: Two techniques. Journal of the Royal Statistical Society. Series D (The Statistician), 21(3):181–194, September 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2986681.

Anonymous:1972:ISM

[143] Anonymous. The Institute of Statisticians minutes of the 22nd Annual General Meeting. Journal of the Royal Statistical Society. Series D (The Statistician), 21(3):195–196, September 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986682.

Neave:1972:SQT

[144] Henry R. Neave. Some quick tests for slippage. Journal of the Royal Statistical Society. Series D (The Statistician), 21 (3):197–208, September 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2986683.

Anonymous:1972:FMc

[145] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 21(3):209, September 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2986677.

Binns:1972:PUA

[146] W. O. Binns. Purpose, understanding and authority in forestry research: Their bearing on statistical consultancy. Journal of the Royal Statistical Society. Series D (The Statistician), 21(3):211– 213, September 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2986684.

Greig-Smith:1972:SPA

[147] P. Greig-Smith. Some problems of analytical data. Journal of the Royal Statistical Society. Series D (The Statistician), 21(3):215–219, September 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2986685.

Huitson:1972:BRBd

[148] Alan Huitson. Book review: Basic Statistics for Business and Economics, by Roger L. Burford. Journal of the Royal Statistical Society. Series D (The Statistician), 21(3):221, September 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2986686.

Lury:1972:BRBd

[149] D. A. Lury. Book review: Facts in Focus, by Central Statistical Office. Journal of the Royal Statistical Society. Series D (The Statistician), 21(3):221– 222, September 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986687.

Ross:1972:BRB

[150] G. J. S. Ross. Book review: Probit Analysis, by D. J. Finney. Journal of the Royal Statistical Society. Series D (The Statistician), 21(3):222, September 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2986688.

Mead:1972:BRB

[151] R. Mead. Book review: Computer Programming in Quantitative Biology, by R. G. Davies. Journal of the Royal Statistical Society. Series D (The Statistician), 21(3):223–224, September 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986689.

Phillips:1972:BRBb

[152] G. D. A. Phillips. Book review: Selected Readings in Econometrics from Econometrica, by J. W. Hooper; M. Nerlove. Journal of the Royal Statistical Society. Series D (The Statistician), 21(3):224– 225, September 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986690.

Harvey:1972:BRB

[153] A. C. Harvey. Book review: Survey Methods in Social Investigation, by C. A. Moser; G. Kalton. Journal of the Royal Statistical Society. Series D (The Statistician), 21(3):225, September 1972. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986691.

Preece:1972:BRBa

[154] D. A. Preece. Book review: Cyclic Designs, by J. A. John; F. W. Wolock; H. A. David; J. M. Cameron; J. A. Speckman. Journal of the Royal Statistical Society. Series D (The Statistician), 21 (3):226, September 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986692.

Basilevsky:1972:BRB

[155] A. Basilevsky. Book review: Factor Analysis as a Statistical Method, by D. N. Lawley; A. E. Maxwell. Journal of the Royal Statistical Society. Series D (The Statistician), 21(3):226– 227, September 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986693.

Jones:1972:BRB

[156] Keith G. Jones. Book review: Understanding Modern Business Mathematics, by A. M. C. Morrison; R. Burden; M. G. Crabtree; Management Planning and Control: Mathematical Models, by John Leslie Livingstone. Journal of the Royal Statistical Society. Series D (The Statistician), 21(3):227–229, September 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986694.

Garside:1972:BRBc

[157] M. J. Garside. Book review: Journal of Statistical Computation and Simulation, by Richard G. Krutchkoff. Journal of the Royal Statistical Society. Series D (The Statistician), 21(3):229, September 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2986696.

Lury:1972:BRBe

[158] D. A. Lury. Book review: Statistical Methods for Research Workers, by R. A. Fisher. Journal of the Royal Statistical Society. Series D (The Statistician), 21 (3):229, September 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986695.

Preece:1972:BRBb

[159] D. A. Preece. Book review: Statistical Design and Analysis of Experiments, by Peter W. M. John. Journal of the Royal Statistical Society. Series D (The Statistician), 21(3):230–232, September 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986697.

Anonymous:1972:BMc

[160] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 21(3):??, September 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2986698.

Anonymous:1972:FMd

[161] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 21(4):i– 290, December 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987122.

Bevan:1972:MIS

[162] J. M. Bevan. Medical information systems: Introduction. Journal of the Royal Statistical Society. Series D (The Statistician), 21(4):233–236, December 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987123.

Bevan:1972:SGR

[163] J. M. Bevan. Some general remarks on the nature and purpose of medical information systems. Journal of the Royal Statistical Society. Series D (The Statistician), 21(4):237–243, December 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987124.

Forsythe:1972:HAA

[164] J. M. Forsythe. Hospital activity analysis. Journal of the Royal Statistical Society. Series D (The Statistician), 21(4): 245–253, December 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987125.

Cust:1972:MUD

[165] George Cust. Morbidity and utilisation data in the local health authority. Journal of the Royal Statistical Society. Series D (The Statistician), 21(4): 255–264, December 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987126.

Ashford:1972:PCG

[166] J. R. Ashford. Patient contacts in general practice in the National Health Service. Journal of the Royal Statistical Society. Series D (The Statistician), 21(4):

265–289, December 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987127.

Davies:1972:MIS

[167] J. G. Davies. Medical information systems for a unified health service: A Scottish view. Journal of the Royal Statistical Society. Series D (The Statistician), 21(4):291–299, December 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987128.

Kalton:1972:RPS

[168] Graham Kalton. The role of population surveys as a source of morbidity and other health data. Journal of the Royal Statistical Society. Series D (The Statistician), 21(4):301–324, December 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987129.

Baldwin:1972:LRH

[169] J. A. Baldwin. Linked record health data systems. Journal of the Royal Statistical Society. Series D (The Statistician), 21 (4):325–338, December 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987130.

Smith:1972:SPO

[170] Alwyn Smith. Summing up: The problems of obtaining information on the health of populations. Journal of the Royal Statistical Society. Series D (The Statistician), 21(4):339– 357, December 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987131.

Anonymous:1972:BMd

[171] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 21(4):??, December 1972. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2987132.

Anonymous:1973:FMa

[172] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 22(1):1–68, March 1973. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2987150.

Moore:1973:MUS

[173] J. Grange Moore. The manager as a user of statistics. Journal of the Royal Statistical Society. Series D (The Statistician), 22(1):3–9, March 1973. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2987151.

Grimshaw:1973:SM

[174] Jack J. Grimshaw. The statistician and management. Journal of the Royal Statistical Society. Series D (The Statistician), 22(1):10–15, March 1973. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987152.

Thomas:1973:MCA

[175] D. A. H. Thomas. Multiple comparisons among means — a review. Journal of the Royal Statistical Society. Series D (The Statistician), 22(1):16–42, March 1973. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987153.

Bissell:1973:MER

[176] A. F. Bissell. Monitoring event rates using varying sample element sizes. Journal of the Royal Statistical Society. Series D (The Statistician), 22(1):43–58, March 1973. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987154.

Walker:1973:NPM

[177] Pamela Walker. Non-parametric methods in quality control. Journal of the Royal Statistical Society. Series D (The Statistician), 22(1):59–67, March 1973. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987155.

Cadman:1973:ASC

[178] M. H. Cadman. Animal spirits or cold calculation? An introduction to investment appraisal. Journal of the Royal Statistical Society. Series D (The Statistician), 22(1):69–79, March 1973. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987156.

Anonymous:1973:BMa

[179] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 22(1):80–88, March 1973. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987158.

Anonymous:1973:ISJ

[180] Anonymous. The Institute of Statisticians June 1972 examination pass list. Journal of the Royal Statistical Society. Series D (The Statistician), 22 (1):81–85, March 1973. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987157.

Anonymous:1973:VI

[181] Anonymous. Volume information. Journal of the Royal Statistical Society. Series D (The Statistician), 22(1):86–87, March 1973. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2987149.

Anonymous:1973:FMb

[182] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 22(2):89–150, June 1973. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2987357.

Williams:1973:E

[183] K. Williams. Editorial. Journal of the Royal Statistical Society. Series D (The Statistician), 22(2):90, June 1973. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987358.

Craddock:1973:FO

[184] J. M. Craddock. Foreword by the organizer. Journal of the Royal Statistical Society. Series D (The Statistician), 22(2):91, June 1973. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987359.

Anonymous:1973:RGS

[185] Anonymous. A report on a geophysical– statistical symposium. Journal of the Royal Statistical Society. Series D (The Statistician), 22(2):92, June 1973. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987360.

Cook:1973:SPG

[186] A. H. Cook. Statistical problems in geophysics. Journal of the Royal Statistical Society. Series D (The Statistician), 22(2):93–102, June 1973. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987361.

Bleasdale:1973:SPP

[187] A. Bleasdale. Statistical problems in practical hydrometeorology. Journal of the Royal Statistical Society. Series D (The Statistician), 22(2):103–123, June 1973. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987362.

Lynden-Bell:1973:QSM

[188] D. Lynden-Bell. Quasar statistics and the most independent variables. Journal of the Royal Statistical Society. Series D (The Statistician), 22(2):125–127, June 1973. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987363.

Kendall:1973:TSA

[189] M. G. Kendall. Techniques in spectral analysis. Journal of the Royal Statistical Society. Series D (The Statistician), 22(2):129–131, June 1973. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987364.

Craddock:1973:PPE

[190] J. M. Craddock. Problems and prospects for eigenvector analysis in meteorology. Journal of the Royal Statistical Society. Series D (The Statistician), 22(2): 133–145, June 1973. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987365.

OConnell:1973:TSA

[191] P. E. O’Connell. Time series analysis. Journal of the Royal Statistical Society. Series D (The Statistician), 22 (2):147–149, June 1973. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987366.

Sprent:1973:FYE

[192] P. Sprent. Frank Yates and experimental design — reflections inspired by his selected papers. Journal of the Royal Statistical Society. Series D (The Statistician), 22(2):151–158, June 1973. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987367.

Anonymous:1973:BMb

[193] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 22(2):159–160, June 1973. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987368.

Anonymous:1973:FMc

[194] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 22(3):161– 164, September 1973. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986765.

Pearson:1973:SHR

[195] E. S. Pearson. Some historical reflections on the introduction of statistical methods in industry. Journal of the Royal Statistical Society. Series D (The Statistician), 22(3):165–179, September 1973. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986766.

Conniffe:1973:CVR

[196] D. Conniffe and Joan Stone. A critical view of ridge regression. Journal of the Royal Statistical Society. Series D (The Statistician), 22(3):181– 187, September 1973. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986767.

Hyett:1973:SAS

[197] G. P. Hyett. Standardizing analysis by social classification. Journal of the Royal Statistical Society. Series D (The Statistician), 22(3):189–198, September 1973. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986768.

Barnett:1973:BDT

[198] V. Barnett. Bayesian, and decision theoretic, methods applied to industrial problems. Journal of the Royal Statistical Society. Series D (The Statistician), 22(3):199–226, September 1973. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986769.

Jenkins:1973:WC

[199] G. M. Jenkins. Written contribution. Journal of the Royal Statistical Society. Series D (The Statistician), 22(3):227– 232, September 1973. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986770.

Craddock:1973:RPG

[200] J. M. Craddock. A reply to Professor G. M. Jenkins. Journal of the Royal Statistical Society. Series D (The Statistician), 22(3):233–235, September 1973. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986771.

Thomas:1973:EMC

[201] D. A. H. Thomas. Errata: Multiple comparison among means. Journal of the Royal Statistical Society. Series D (The Statistician), 22(3):236, September 1973. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2986772.

Anonymous:1973:BMc

[202] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 22(3):237– 240, September 1973. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986773.

Liddell:1973:MVR

[203] Douglas Liddell. Multi-variate response in the matched (one sample) situation. Journal of the Royal Statistical Society. Series D (The Statistician), 22(4): 241–255, December 1973. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986822.

VanGerven:1973:UMS

[204] Dennis P. Van Gerven and Gail B. Oakland. Univariate and multivariate statistical model in the analysis of human sexual dimorphism. Journal of the Royal Statistical Society. Series D (The Statistician), 22(4):256–268, December 1973. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986823.

Neave:1973:PSS

[205] Henry R. Neave. A power study of some tests for slippage. Journal of the Royal Statistical Society. Series D (The Statistician), 22(4):269–280, December 1973. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986824.

Chatfield:1973:MRS

[206] C. Chatfield and C. M. Theobald. Mixtures and random sums. Journal of the Royal Statistical Society. Series D (The Statistician), 22(4):281–287, December 1973. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2986825.

Murphy:1973:AME

[207] B. P. Murphy and C. P. Chalmers. Assessing modern electronic calculators.

Journal of the Royal Statistical Society. Series D (The Statistician), 22(4): 289–297, December 1973. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986826.

Jeffers:1973:BSG

[208] J. N. R. Jeffers. A basic subroutine for Geary’s contiguity ratio. Journal of the Royal Statistical Society. Series D (The Statistician), 22(4):299– 302, December 1973. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986827.

Fox:1973:BRB

[209] R. A. Fox. Book review: Practical Business Statistics, by F. E. Croxton; D. J. Cowden; B. W. Bolch. Journal of the Royal Statistical Society. Series D (The Statistician), 22(4):303, December 1973. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986828.

Davis:1973:BRB

[210] E. J. Davis. Book review: Scientific Marketing Research, by Gerald Albaum; M. Venkatesan. Journal of the Royal Statistical Society. Series D (The Statistician), 22(4):303–304, December 1973. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986829.

Fuller:1973:BRB

[211] M. F. Fuller. Book review: Practical Nonparametric Statistics, by W. J. Conover; Nonparametric Statistical Inference, by Jean Dickinson Gibbons. Journal of the Royal Statistical Society. Series D (The Statistician), 22(4):

304–305, December 1973. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986830.

Jeffers:1973:BRB

[212] J. N. R. Jeffers. Book review: Multivariate Morphometrics, by R. E. Blackith; R. A. Reyment; Mathematical Taxanomy, by Nicholas Jardine; Robin Sibson. Journal of the Royal Statistical Society. Series D (The Statistician), 22(4): 305–306, December 1973. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986831.

Andrew:1973:BRB

[213] Ian A. Andrew. Book review: Proceeding of the Third Conference of the Advisory Group of Forest Statisticians, Jouy-en-Josas, 1970, by M. R. Tomassone. Journal of the Royal Statistical Society. Series D (The Statistician), 22(4): 306–307, December 1973. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986832.

Anonymous:1973:BMd

[214] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 22(4):309– 312, December 1973. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986833.

Anonymous:1973:FMd

[215] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 22(4):??, December 1973. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2986821.

Anonymous:1974:FMa

[216] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 23(1):i–iv, March 1974. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2987259.

Jones:1974:SIA

[217] Keith G. Jones. Statistical investment analysis and accounting: An introduction. Journal of the Royal Statistical Society. Series D (The Statistician), 23 (1):1–3, March 1974. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987260.

Dickenson:1974:SSA

[218] J. P. Dickenson. Some statistical aspects of portfolio analysis. Journal of the Royal Statistical Society. Series D (The Statistician), 23(1):5–16, March 1974. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987261.

Russell:1974:EST

[219] A. H. Russell. Estimation of B in the Sharpe/Tobin capital asset evaluation model. Journal of the Royal Statistical Society. Series D (The Statistician), 23 (1):17–30, March 1974. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987262.

Hobbs:1974:DPS

[220] David H. S. Hobbs. The development of a practical share evaluation model.

Journal of the Royal Statistical Society. Series D (The Statistician), 23 (1):31–56, March 1974. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987263.

McRae:1974:PCE

[221] T. W. McRae. The problem of containing error in highly skewed accounting populations. Journal of the Royal Statistical Society. Series D (The Statistician), 23(1):57–65, March 1974. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987264.

Symon:1974:PSE

[222] Iain W. Symon. The place of statistics in the education of chartered accountants. Journal of the Royal Statistical Society. Series D (The Statistician), 23 (1):67–70, March 1974. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987265.

Sully:1974:SSA

[223] J. M. Sully. Statistical sampling in auditing. Journal of the Royal Statistical Society. Series D (The Statistician), 23 (1):71–80, March 1974. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987266.

Anonymous:1974:VI

[224] Anonymous. Volume information. Journal of the Royal Statistical Society. Series D (The Statistician), 23(1):81, March 1974. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987258.

Anonymous:1974:BMa

[225] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 23(1):82–84, March 1974. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2987267.

Anonymous:1974:ODC

[226] Anonymous. Obituary: Dr. C. Oswald George. Journal of the Royal Statistical Society. Series D (The Statistician), 23(2):85–88, June 1974. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988007.

Hammersley:1974:ST

[227] J. M. Hammersley. Statistical tools. Journal of the Royal Statistical Society. Series D (The Statistician), 23 (2):89–106, June 1974. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988008.

McKenzie:1974:CSF

[228] Ed McKenzie. A comparison of standard forecasting systems with the Box– Jenkins approach. Journal of the Royal Statistical Society. Series D (The Statistician), 23(2):107–116, June 1974. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988009.

Button:1974:CRT

[229] K. J. Button. A critical review of techniques used for forecasting car ownership in discrete areas. Journal of the Royal Statistical Society. Series D (The Statistician), 23(2):117–128, June 1974.

CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988010.

Holmes:1974:GMA

[230] D. I. Holmes. Graphical methods for the analysis of discrete data. Journal of the Royal Statistical Society. Series D (The Statistician), 23(2):129–134, June 1974. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988011.

Bartholomew:1974:BRB

[231] D. J. Bartholomew. Book review: Scientific Truth and Statistical Method, by M. Boldrini. Journal of the Royal Statistical Society. Series D (The Statistician), 23(2):135–136, June 1974. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988012.

Grimshaw:1974:BRB

[232] J. J. Grimshaw. Book review: The Statistician and the Manager, by Ken Williams. Journal of the Royal Statistical Society. Series D (The Statistician), 23(2):136, June 1974. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2988013.

Lancaster:1974:BRB

[233] T. Lancaster. Book review: Frontiers of Quantitative Economics, by M. D. Intriligator. Journal of the Royal Statistical Society. Series D (The Statistician), 23(2):136–137, June 1974. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988014.

Hill:1974:BRB

[234] Roger Hill. Book review: The Growth and Role of U.K. Financial Institutions 1880–1962, by David K. Sheppard. Journal of the Royal Statistical Society. Series D (The Statistician), 23 (2):137–138, June 1974. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988015.

Harvey:1974:BRB

[235] A. C. Harvey. Book review: The Statistical Analysis of Time Series, by T. W. Anderson. Journal of the Royal Statistical Society. Series D (The Statistician), 23(2):138–140, June 1974. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988016.

Spratt:1974:BRB

[236] E. B. Spratt. Book review: Computation and Communication, by G. Emery. Journal of the Royal Statistical Society. Series D (The Statistician), 23(2):140, June 1974. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988017.

Garside:1974:BRB

[237] M. J. Garside. Book review: More Fortran Programs for Economists, by Lucy J. Slater. Journal of the Royal Statistical Society. Series D (The Statistician), 23(2):140–141, June 1974. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988018.

Jones:1974:BRB

[238] Keith G. Jones. Book review: Lecture and Problem Materials in Quantitative Methods, by Charles L. Bostwick; Dennis E. Grawoig; Edward G. Rodgers; William J. Thampson Evanston. Journal of the Royal Statistical Society. Series D (The Statistician), 23(2):141, June 1974. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2988019.

Phillips:1974:BRB

[239] Lawrence D. Phillips. Book review: Empirical Bayes Methods, by J. S. Maritz. Journal of the Royal Statistical Society. Series D (The Statistician), 23(2): 141–142, June 1974. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988020.

John:1974:BRB

[240] J. A. John. Book review: Mathematical Statistics, by John E. Freund. Journal of the Royal Statistical Society. Series D (The Statistician), 23(2): 142–143, June 1974. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988021.

Holgate:1974:BRB

[241] P. Holgate. Book review: The Analysis of Biological Population, by Mark Williamson. Journal of the Royal Statistical Society. Series D (The Statistician), 23(2):143–144, June 1974. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988022.

Preece:1974:BRB

[242] D. A. Preece. Book review: BIOMETRIKA Tables for Statisticians (Volume 2), by E. S. Pearson; H. O. Hartley. Journal of the Royal Statistical Society. Series D (The Statistician), 23(2):144–145, June 1974. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2988023.

Anonymous:1974:BMb

[243] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 23(2):146–148, June 1974. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2988024.

Anonymous:1974:FMb

[244] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 23(2):??, June 1974. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988006.

Taylor:1974:GSI

[245] Peter J. Taylor and John Goddard. Geography and statistics: An introduction. Journal of the Royal Statistical Society. Series D (The Statistician), 23(3/4):149–155, September/December 1974. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2987578.

Anonymous:1974:FMc

[246] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 23(3/4):156, September/December 1974. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987577.

Gudgin:1974:PGR

[247] G. Gudgin and J. B. Thornes. Probability in geographic research: Applications and problems. Journal of the Royal Statistical Society. Series D (The Statistician), 23(3/4):157–177, September/December 1974. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987579.

Collins:1974:MMG

[248] Lyndhurst Collins, Roy Drewett, and Rob Ferguson. Markov models in geography. Journal of the Royal Statistical Society. Series D (The Statistician), 23(3/4):179–209, September/December 1974. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987580.

Unwin:1974:SAS

[249] D. J. Unwin and L. W. Hepple. The statistical analysis of spatial series. Journal of the Royal Statistical Society. Series D (The Statistician), 23(3/4):211– 227, September/December 1974. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987581.

Wilson:1974:PSS

[250] A. G. Wilson and P. H. Rees. Population statistics and spatial demographic accounts. Journal of the Royal Statistical Society. Series D (The Statistician), 23(3/4):229–257, September/December 1974. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2987582.

Clark:1974:AFA

[251] D. Clark, W. K. D. Davies, and R. J. Johnston. The application of factor analysis in human geography. Journal of the Royal Statistical Society. Series D (The Statistician), 23(3/4):259–281, September/December 1974. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2987583.

Mather:1974:MMG

[252] P. Mather and S. Openshaw. Multivariate methods and geographical data. Journal of the Royal Statistical Society. Series D (The Statistician), 23(3/4): 283–308, September/December 1974. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987584.

Anonymous:1974:BMc

[253] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 23(3/4):309–311, September/December 1974. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2987585.

Harvey:1975:SAE

[254] A. C. Harvey. Spectral analysis in economics. Journal of the Royal Statistical Society. Series D (The Statistician), 24(1):1–36, March 1975. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988260.

Everitt:1975:VTR

[255] B. S. Everitt and P. Nicholls. Visual techniques for representing multivariate data. Journal of the Royal Statistical Society. Series D (The Statistician), 24 (1):37–49, March 1975. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988261.

Hyett:1975:FRT

[256] G. P. Hyett and J. R. McKenzie. Forecasting of residential telephone penetration 1974–81 by use of the technique of social sections. Journal of the Royal Statistical Society. Series D (The Statistician), 24(1):51–60, March 1975. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988262.

Smith:1975:RRS

[257] A. F. M. Smith, M. Goldstein, D. Conniffe, and Joan Stone. Ridge regression: Some comments on a paper of conniffe and stone. Journal of the Royal Statistical Society. Series D (The Statistician), 24(1):61–68, March 1975. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988263.

Murphy:1975:EUW

[258] Allan H. Murphy. Expressing the uncertainty in weather forecasts. Journal of the Royal Statistical Society. Series D (The Statistician), 24(1):69–71, March 1975. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2988264.

Anonymous:1975:VI

[259] Anonymous. Volume information. Journal of the Royal Statistical Society. Series D (The Statistician), 24(1):73, March 1975. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2988258.

Anonymous:1975:BMa

[260] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 24(1):74–78, March 1975. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2988265.

Anonymous:1975:FMa

[261] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 24(1):??, March 1975. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2988259.

Bissell:1975:JTT

[262] A. F. Bissell and R. A. Ferguson. The jackknife — toy, tool or two-edged weapon? Journal of the Royal Statistical Society. Series D (The Statistician), 24(2):79–100, June 1975. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987663.

Swee-Hock:1975:MHK

[263] Saw Swee-Hock. Methodology of the Hong Kong Consumer Price Index 1963– 1973. Journal of the Royal Statistical Society. Series D (The Statistician), 24 (2):101–114, June 1975. CODEN ????

ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987664.

Furlong:1975:TPD

[264] J. D. G. Furlong and R. Szreter. The trend of the performance differential between leading men and women athletes, 1951–67. Journal of the Royal Statistical Society. Series D (The Statistician), 24(2):115–128, June 1975. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987665.

Phipps:1975:LCF

[265] D. F. Phipps. The logistic curve: A fitting technique. Journal of the Royal Statistical Society. Series D (The Statistician), 24(2):129–136, June 1975. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987666.

Moroney:1975:DON

[266] Davitt Moroney. Divinity in odd numbers. Journal of the Royal Statistical Society. Series D (The Statistician), 24 (2):137–143, June 1975. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987667.

Anonymous:1975:BMb

[267] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 24(2):144–156, June 1975. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987678.

Reuter:1975:BRB

[268] G. E. H. Reuter. Book review: Regenerative Phenomena, by J. F. C. Kingman. Journal of the Royal Statistical Society. Series D (The Statistician), 24 (2):145–146, June 1975. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987668.

Rowcroft:1975:BRB

[269] John E. Rowcroft. Book review: Distributions in Statistics: Continuous Multivariate Distributions, by N. L. Johnson; S. Kotz. Journal of the Royal Statistical Society. Series D (The Statistician), 24(2):146–147, June 1975. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987669.

Hiorns:1975:BRB

[270] R. W. Hiorns. Book review: Biomathematics Volume 1 — Mathematical Topics in Population Genetics, by Ken-Ichi Kojima. Journal of the Royal Statistical Society. Series D (The Statistician), 24(2):147–148, June 1975. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987670.

Grimshaw:1975:BRBa

[271] J. J. Grimshaw. Book review: Statistical Analysis for Business: A Conceptual Approach, by Lee H. Smith; Donald R. Williams. Journal of the Royal Statistical Society. Series D (The Statistician), 24(2):148–149, June 1975. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987671.

Grimshaw:1975:BRBb

[272] J. J. Grimshaw. Book review: Modern Statistics for Business and Economics, by W. L. L’Esperance. Journal of the Royal Statistical Society. Series D (The Statistician), 24(2):149, June 1975. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987672.

Wilson:1975:BRB

[273] Eveline Wilson. Book review: FOR-TRAN Techniques with Special Reference to Non-Numerical Applications, by A. Colin Day. Journal of the Royal Statistical Society. Series D (The Statistician), 24(2):149–150, June 1975. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987673.

Lury:1975:BRB

[274] D. A. Lury. Book review: Advances in Social Research, by R. Cochrane. Journal of the Royal Statistical Society. Series D (The Statistician), 24(2): 150–151, June 1975. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987674.

Pearce:1975:BRB

[275] S. C. Pearce. Book review: Collected Papers of R. A. Fisher. (Volume 2), by R. A. Fisher. Journal of the Royal Statistical Society. Series D (The Statistician), 24(2):151, June 1975. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987675.

Grimshaw:1975:BRBc

[276] J. J. Grimshaw. Book review: Statistics in Theory and Practice, by L. R. Connor; A. J. H. Morrell. Journal of the Royal Statistical Society. Series D (The Statistician), 24(2):152, June 1975. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987677.

Preece:1975:BRBa

[277] D. A. Preece. Book review: Statistics for Biologists, by R. C. Campbell. Journal of the Royal Statistical Society. Series D (The Statistician), 24(2):152, June 1975. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987676.

Anonymous:1975:FMb

[278] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 24(2):??, June 1975. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987662.

Anonymous:1975:E

[279] Anonymous. Editorial. Journal of the Royal Statistical Society. Series D (The Statistician), 24(3):157, September 1975. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2987776.

Anonymous:1975:FMc

[280] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 24(3):158, September 1975. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987775.

Hammersley:1975:WOS

[281] J. M. Hammersley. The wide open spaces. Journal of the Royal Statistical Society. Series D (The Statistician), 24 (3):159–160, September 1975. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987777.

Moran:1975:USG

[282] P. A. P. Moran. The use of statistics in geography. Journal of the Royal Statistical Society. Series D (The Statistician), 24(3):161–162, September 1975. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987778.

Benjamin:1975:DV

[283] Bernard Benjamin. A demographic viewpoint. Journal of the Royal Statistical Society. Series D (The Statistician), 24(3):163–165, September 1975. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987779.

Kingman:1975:MMS

[284] J. F. C. Kingman. Markov models for spatial variation. Journal of the Royal Statistical Society. Series D (The Statistician), 24(3):167–174, September 1975. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987780.

Bartholomew:1975:CMM

[285] D. J. Bartholomew. Comments on Markov models in geography. Journal of the Royal Statistical Society. Series D (The Statistician), 24(3):175– 177, September 1975. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987781.

Besag:1975:SAN

[286] Julian Besag. Statistical analysis of nonlattice data. Journal of the Royal Statistical Society. Series D (The Statistician), 24(3):179–195, September 1975. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987782.

Granger:1975:AAI

[287] C. W. J. Granger. Aspects of the analysis and interpretation of temporal and spatial data. Journal of the Royal Statistical Society. Series D (The Statistician), 24(3):197–210, September 1975. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987783.

Scott:1975:MAG

[288] J. F. Scott. Multivariate analysis in geography: Some comments. Journal of the Royal Statistical Society. Series D (The Statistician), 24(3):211– 216, September 1975. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987784.

Grimshaw:1975:SMP

[289] J. J. Grimshaw. Statistical and management problems in clinical trials. Journal of the Royal Statistical Society. Series D (The Statistician), 24(3):217– 226, September 1975. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987785.

Grimshaw:1975:BRBd

[290] J. J. Grimshaw. Book review: Modern Mathematics for Managers, by Ken Williams. Journal of the Royal Statistical Society. Series D (The Statistician), 24(3):227, September 1975. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987786.

Benjamin:1975:BRB

[291] Bernard Benjamin. Book review:

Sources of Social Statistics, by Bernard Edwards. Journal of the Royal Statistical Society. Series D (The Statistician), 24(3):227–228, September 1975. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987787.

Ilersic:1975:BRB

[292] A. R. Ilersic. Book review: Trends in British Society Since 1900, by A. H. Halsey. Journal of the Royal Statistical Society. Series D (The Statistician), 24 (3):228–230, September 1975. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987788.

Preece:1975:BRBb

[293] D. A. Preece. Book review: Tables of Two-Associate-Class Partially Balanced Designs, by Willard H. Clatworthy. Journal of the Royal Statistical Society. Series D (The Statistician), 24(3): 230–231, September 1975. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987789.

Marchant:1975:BRB

[294] L. Marchant. Book review: Consumer Market Research Handbook, by Robert M. Worcester. Journal of the Royal Statistical Society. Series D (The Statistician), 24(3):231, September 1975. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987790.

Phipps:1975:ALC

[295] D. F. Phipps. An amendment: The logistic curve: A fitting technique. Journal of the Royal Statistical Society. Series D (The Statistician), 24(3):232– 233, September 1975. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987791.

Anonymous:1975:BMc

[296] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 24(3):234, September 1975. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2987792.

Gerson:1975:TUP

[297] Marion Gerson. The techniques and uses of probability plotting. Journal of the Royal Statistical Society. Series D (The Statistician), 24(4):235– 257, December 1975. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987921.

Conniffe:1975:DSR

[298] D. Conniffe. Double sampling with regression — extension to the case of unequal regression coefficients. Journal of the Royal Statistical Society. Series D (The Statistician), 24(4):259– 266, December 1975. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987922.

Hsiang:1975:BVR

[299] T. C. Hsiang. A Bayesian view on ridge regression. Journal of the Royal Statistical Society. Series D (The Statistician), 24(4):267–268, December 1975. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987923.

Evans:1975:SRE

[300] Graham J. Evans and Colin J. Roberts.

Short-run employment functions in simulation and forecasting. Journal of the Royal Statistical Society. Series D (The Statistician), 24(4):269–282, December 1975. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987924.

Anderson:1975:MAP

[301] O. D. Anderson. Moving average processes. Journal of the Royal Statistical Society. Series D (The Statistician), 24 (4):283–297, December 1975. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987925.

Wood:1975:MMA

[302] P. D. P. Wood. A method of multivariate analysis by order statistics. Journal of the Royal Statistical Society. Series D (The Statistician), 24(4):299– 305, December 1975. CODEN ????

ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987926.

Smith:1975:BRB

[303] A. F. M. Smith. Book review: Contiguity of Probability Measures: Some Applications in Statistics, by George G. Roussas. Journal of the Royal Statistical Society. Series D (The Statistician), 24(4): 307–308, December 1975. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987927.

Prescott:1975:BRB

[304] P. Prescott. Book review: Robust Estimates of Location: Survey and Advances, by D. R. Andrews; P. J. Bickel; F. R. Hampei; P. J. Huber; W. H. Rogers; J. W. Tukey. Journal of the Royal Statistical Society. Series D (The Statistician), 24(4):308–309, December 1975. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2987928.

Harvey:1975:BRB

[305] A. C. Harvey. Book review: Regression Estimation from Grouped Observations, by Y. Haitovsky. Journal of the Royal Statistical Society. Series D (The Statistician), 24(4):309–310, December 1975. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987929.

Gordon:1975:BRB

[306] Ian Gordon. Book review: N.I.E.S.R. Regional Papers II: Regional Unemployment Differences in Great Britain, and Interregional Migration Models and Their Application to Great Britain, by P. C. Cheshire; R. Weeden. Journal of the Royal Statistical Society. Series D (The Statistician), 24(4):310, December 1975. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987930.

Anonymous:1975:BMd

[307] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 24(4):311– 312, December 1975. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987931.

Anonymous:1975:FMd

[308] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 24(4):??, December 1975. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987920.

Anderson:1976:FW

[309] O. D. Anderson. The first word.

Journal of the Royal Statistical Society. Series D (The Statistician), 25 (1):1–2, March 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988126.

McKenzie:1976:CSS

[310] Ed McKenzie. A comparison of some standard seasonal forecasting systems. Journal of the Royal Statistical Society. Series D (The Statistician), 25 (1):3–14, March 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988127.

Edwards:1976:FP

[311] A. W. F. Edwards. Fiducial probability. Journal of the Royal Statistical Society. Series D (The Statistician), 25 (1):15–35, March 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988128.

Manly:1976:EDT

[312] B. F. J. Manly. Exponential data transformations. Journal of the Royal Statistical Society. Series D (The Statistician), 25(1):37–42, March 1976. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988129.

Kanji:1976:SPA

[313] G. K. Kanji. Sensitivity of the power in analysis of variance to the departures from the in-built assumptions. Journal of the Royal Statistical Society. Series D (The Statistician), 25 (1):43–48, March 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988130.

Williams:1976:FPG

[314] Ken Williams. The failure of Pearson’s goodness of fit statistic. Journal of the Royal Statistical Society. Series D (The Statistician), 25(1):49, March 1976. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988131.

Williams:1976:AMC

[315] Ken Williams. Analysis of multidimensional contingency tables. Journal of the Royal Statistical Society. Series D (The Statistician), 25(1):51–58, March 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988132.

Wilson:1976:SDS

[316] Alan Wilson and Philip Rees. Spatial demography: Some comments. Journal of the Royal Statistical Society. Series D (The Statistician), 25(1):59–61, March 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2988133.

Anderson:1976:CMA

[317] O. D. Anderson. Corrections: Moving average processes. Journal of the Royal Statistical Society. Series D (The Statistician), 25(1):62, March 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2988135.

Murphy:1976:CEU

[318] Allan H. Murphy. Corrections: Expressing the uncertainty in weather forecasts. Journal of the Royal Statistical Society. Series D (The Statistician), 25(1):62, March 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988134.

Kendall:1976:BRB

[319] M. G. Kendall. Book review: Clustering Algorithms, by J. A. Hartigan. Journal of the Royal Statistical Society. Series D (The Statistician), 25 (1):63, March 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988136.

Anderson:1976:BRBa

[320] O. D. Anderson. Book review: Sequential Methods in Statistics, by G. Barrie Wetherill. Journal of the Royal Statistical Society. Series D (The Statistician), 25(1):63–64, March 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988137.

Anderson:1976:BRBb

[321] O. D. Anderson. Book review: Statistical Inference, by S. D. Silvey. Journal of the Royal Statistical Society. Series D (The Statistician), 25(1):64, March 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2988138.

Mead:1976:BRB

[322] R. Mead. Book review: Linear Statistical Inference and its Statistical Applications, by C. R. Rao. Journal of the Royal Statistical Society. Series D (The Statistician), 25(1):64–65, March 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988139.

Harvey:1976:BRBa

[323] A. C. Harvey. Book review: Time Series, by M. G. Kendall. Journal of the Royal Statistical Society. Series D (The Statistician), 25(1):65–66, March 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988140.

Bartholomew:1976:BRBa

[324] D. J. Bartholomew. Book review: An Objective Theory of Probability, by D.

A. Gillies. Journal of the Royal Statistical Society. Series D (The Statistician), 25(1):66–68, March 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988141.

Bartholomew:1976:BRBb

[325] D. J. Bartholomew. Book review: Scientific Inference, 3rd Ed, by Harold Jeffreys. Journal of the Royal Statistical Society. Series D (The Statistician), 25 (1):68–69, March 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988142.

Chapas:1976:BRB

[326] L. C. Chapas. Book review: Introduction to Biostatistics, by Robert R. Sokal; F. James Rohlf; W. H. Freeman. Journal of the Royal Statistical Society. Series D (The Statistician), 25 (1):69–70, March 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988143.

Marshall:1976:BRB

[327] T. F. de C. Marshall. Book review: Statistical Methods for Rates and Proportions, by J. L. Fleiss. Journal of the Royal Statistical Society. Series D (The Statistician), 25(1):70–71, March 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988144.

Cooper:1976:BRB

[328] R. A. Cooper. Book review: Econometric Methods, 2nd Ed., by J. Johnston; Principles of Econometrics, by H. Theil. Journal of the Royal Statistical Society. Series D (The Statistician), 25 (1):71–72, March 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988145.

Phillips:1976:BRB

[329] G. D. A. Phillips. Book review: Optimisation Under Constraints, by P. Whittle. Journal of the Royal Statistical Society. Series D (The Statistician), 25 (1):72, March 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988146.

Bartholomew:1976:BRBc

[330] D. J. Bartholomew. Book review: Probability and Statistical Inference in Ancient and Medieval Jewish Literature, by Nachum L. Rabinovitch. Journal of the Royal Statistical Society. Series D (The Statistician), 25(1):73, March 1976. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988147.

Davies:1976:BRB

[331] G. S. Davies. Book review: Stochastic Models for Social Processes, by D. J. Bartholomew. Journal of the Royal Statistical Society. Series D (The Statistician), 25(1):73–74, March 1976. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988148.

Liddell:1976:BRB

[332] F. D. K. Liddell. Book review: Cancer in the United States, by Abraham M. Lilienfield; Morton I. Levin; Irving I. Kessler. Journal of the Royal Statistical Society. Series D (The Statistician), 25(1):74–75, March 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988149.

Mardia:1976:BRB

[333] K. V. Mardia. Book review: Characterization Problems in Mathematical Statistics, by A. M. Kagan; Yu. V. Linnik; C. Radhakrishna Rao; B. Ramachandran. Journal of the Royal Statistical Society. Series D (The Statistician), 25 (1):75–76, March 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988150.

Anonymous:1976:VIa

[334] Anonymous. Volume information. Journal of the Royal Statistical Society. Series D (The Statistician), 25(1):77–78, March 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2988124.

Anonymous:1976:BMa

[335] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 25(1):??, March 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2988151.

Anonymous:1976:FMa

[336] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 25(1):??, March 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2988125.

Anderson:1976:STH

[337] O. D. Anderson. Statistics teaching in higher education. Journal of the Royal Statistical Society. Series D (The Statistician), 25(2):79–94, June 1976. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987641.

Scott:1976:PPT

[338] J. F. Scott. Practical projects in the teaching of statistics at universities. Journal of the Royal Statistical Society. Series D (The Statistician), 25 (2):95–108, June 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987642.

Pike:1976:SGT

[339] Derek J. Pike. Statistical games as teaching aids. Journal of the Royal Statistical Society. Series D (The Statistician), 25(2):109–115, June 1976. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987643.

Griffiths:1976:PWS

[340] J. D. Griffiths and B. E. Evans. Project work in statistics. Journal of the Royal Statistical Society. Series D (The Statistician), 25(2):117–123, June 1976. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987644.

Griffiths:1976:PTP

[341] J. D. Griffiths and B. E. Evans. Practical training periods for statisticians. Journal of the Royal Statistical Society. Series D (The Statistician), 25(2):

125–128, June 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987645.

Braithwaite:1976:CDR

[342] G. R. Braithwaite. Coursework, a discussion of its role and assessment. Journal of the Royal Statistical Society. Series D (The Statistician), 25(2): 129–135, June 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987646.

Gilchrist:1976:ES

[343] Warren Gilchrist. Examining statistics. Journal of the Royal Statistical Society. Series D (The Statistician), 25 (2):137–145, June 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987647.

Anonymous:1976:FMb

[344] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 25(2):146, June 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987640.

Moore:1976:TSB

[345] Peter G. Moore. The teaching of statistics at a business school. Journal of the Royal Statistical Society. Series D (The Statistician), 25(2):147–154, June 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987648.

Rosenbaum:1976:CSC

[346] S. Rosenbaum. The Civil Service College. Journal of the Royal Statistical Society. Series D (The Statistician), 25 (2):155–167, June 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987649.

Kanji:1976:CSP

[347] G. K. Kanji. Correction: Sensitivity of the power in analysis of variance to the departures from the in-built assumptions. Journal of the Royal Statistical Society. Series D (The Statistician), 25(2):167, June 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987650.

Anonymous:1976:BMb

[348] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 25(2):168–172, June 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987651.

Anderson:1976:SSE

[349] O. D. Anderson. The statistician, statistical education and the Institute. Journal of the Royal Statistical Society. Series D (The Statistician), 25(3):173– 178, September 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987831.

Kerridge:1976:MM

[350] D. F. Kerridge. The menace of mathematics. Journal of the Royal Statistical Society. Series D (The Statistician), 25 (3):179–189, September 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987832.

Anonymous:1976:FMc

[351] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 25(3):190– 212, September 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987830.

Ehrenberg:1976:SC

[352] A. S. C. Ehrenberg. Symbols or concepts? Journal of the Royal Statistical Society. Series D (The Statistician), 25 (3):191–193, September 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987833.

Ehrenberg:1976:WMP

[353] A. S. C. Ehrenberg. We must preach what is practised: A radical review of statistical teaching. Journal of the Royal Statistical Society. Series D (The Statistician), 25(3):195–208, September 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987834.

Holmes:1976:TDA

[354] D. I. Holmes. Teaching data analysis to first year students at a polytechnic. Journal of the Royal Statistical Society. Series D (The Statistician), 25(3):209– 211, September 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987835.

Mead:1976:SCU

[355] R. Mead. Statistical consulting in a university. Journal of the Royal Statistical Society. Series D (The Statistician), 25 (3):213–218, September 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987836.

Farid:1976:EAS

[356] S. M. Farid. On estimating age-specific nuptiality rates. Journal of the Royal Statistical Society. Series D (The Statistician), 25(3):219–224, September 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987837.

Wynn:1976:UEU

[357] H. P. Wynn. An unbiased estimator under SRS of a small change in a proportion. Journal of the Royal Statistical Society. Series D (The Statistician), 25(3): 225–228, September 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987838.

Anderson:1976:BRBc

[358] O. D. Anderson. Book review: Multivariate Analysis, by Maurice Kendall. Journal of the Royal Statistical Society. Series D (The Statistician), 25(3): 229, September 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987839.

Lury:1976:BRB

[359] D. A. Lury. Book review: Statistics Made Relevant: A Casebook of Real Life Examples, by P. Baum; E. M. Schauer. Journal of the Royal Statistical Society. Series D (The Statistician), 25(3):229– 230, September 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987840.

Baldwin:1976:BRB

[360] M. Baldwin. Book review: An Introduction to Applied Macro-Economics, by E. Kuh; R. L. Schmalensee. Journal of the Royal Statistical Society. Series D (The Statistician), 25(3):230– 231, September 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987841.

Morgan:1976:BRB

[361] B. J. T. Morgan. Book review: A Primer of Signal Detection Theory, by D. Mc-Nicol; Table of d′ and β, by P. R. Freeman. Journal of the Royal Statistical Society. Series D (The Statistician), 25(3): 231–233, September 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987842.

MacDonald:1976:BRB

[362] R. R. MacDonald. Book review: Likelihood, by A. W. F. Edwards. Journal of the Royal Statistical Society. Series D (The Statistician), 25(3):233– 235, September 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987843.

Harvey:1976:BRBb

[363] A. C. Harvey. Book review: Distributed Lags: Problems of Estimation and Formulation, by P. J. Dhrymes. Journal of the Royal Statistical Society. Series D (The Statistician), 25(3):235– 236, September 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987844.

Huitson:1976:BRB

[364] A. Huitson. Book review: Applied Statistics: Analysis of Variance and Regression, by Olive Jean Dunn; Virginia A. Clark. Journal of the Royal Statistical Society. Series D (The Statistician), 25 (3):236, September 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987845.

Anonymous:1976:BMc

[365] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 25(3):??, September 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2987846.

Anderson:1976:LSS

[366] O. D. Anderson. The last of seventysix. Journal of the Royal Statistical Society. Series D (The Statistician), 25(4): 237–246, December 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988081.

Pollock:1976:SST

[367] K. H. Pollock and I. M. Wilson. Statistics service teaching in universities. Journal of the Royal Statistical Society. Series D (The Statistician), 25(4): 247–252, December 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988082.

Pollock:1976:BMC

[368] K. H. Pollock. Building models of capture-recapture experiments. Journal of the Royal Statistical Society. Series D (The Statistician), 25(4):253– 259, December 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988083.

Anonymous:1976:JET

[369] Anonymous. June 1976 examination:

Top three passes. Journal of the Royal Statistical Society. Series D (The Statistician), 25(4):260, December 1976. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988084.

Barnett:1976:SJA

[370] V. Barnett. The statistician: Jack of all trades, master of one? Journal of the Royal Statistical Society. Series D (The Statistician), 25(4):261– 279, December 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988085.

Anonymous:1976:FMd

[371] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 25(4):280, December 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2988080.

McClean:1976:CLT

[372] Sally McClean. A comparison of the lognormal and transition models of wastage. Journal of the Royal Statistical Society. Series D (The Statistician), 25(4):281–294, December 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2988086.

Liddell:1976:PTC

[373] Douglas Liddell. Practical tests of 2 × 2 contingency tables. Journal of the Royal Statistical Society. Series D (The Statistician), 25(4):295–304, December 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988087.

Steece:1976:ICV

[374] Bert M. Steece. The implications of a certain variable selection criterion in regression analysis. Journal of the Royal Statistical Society. Series D (The Statistician), 25(4):305–307, December 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988088.

Whitacre:1976:BRB

[375] R. J. Whitacre. Book review: Problems in Linear and Non-Linear Programming, by S. Vajda. Journal of the Royal Statistical Society. Series D (The Statistician), 25(4):307, December 1976. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988089.

Anderson:1976:BRBd

[376] O. D. Anderson. Book review: The Analysis of Time Series: Theory and Practice, by C. Chatfield. Journal of the Royal Statistical Society. Series D (The Statistician), 25(4):308, December 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988090.

Anderson:1976:BRBe

[377] O. D. Anderson. Book review: Time-Series, 2nd edn, by Maurice Kendall.

Journal of the Royal Statistical Society. Series D (The Statistician), 25(4): 308–310, December 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988091.

Heuts:1976:BRB

[378] R. M. J. Heuts. Book review: Time Series Analysis and Forecasting: The Box-Jenkins Approach, by O. D. Anderson. Journal of the Royal Statistical Society. Series D (The Statistician), 25(4): 310–311, December 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988092.

Anderson:1976:BRBf

[379] O. D. Anderson. Book review: The Advanced Theory of Statistics, Vol. 3: Design and Analysis, and Time- Series. 3rd Edn, by Maurice Kendall; Alan Stuart. Journal of the Royal Statistical Society. Series D (The Statistician), 25(4): 311–313, December 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988093.

Preece:1976:BRB

[380] D. A. Preece. Book review: Statistics and Society: Data Collection and Interpretation, by W. T. Federer. Journal of the Royal Statistical Society. Series D (The Statistician), 25(4):314– 315, December 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988094.

Anonymous:1976:VIb

[381] Anonymous. Volume information. Journal of the Royal Statistical Society. Series D (The Statistician), 25(4):315– 316, December 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988079.

Anonymous:1976:BMd

[382] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 25(4):??, December 1976. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2988095.

Liddell:1977:MRM

[383] Douglas Liddell. Multivariate response in more than one sample. Journal of the Royal Statistical Society. Series D (The Statistician), 26(1):1–15, March 1977. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988215.

Anonymous:1977:FMa

[384] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 26(1):16–45, March 1977. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2988214.

Fielding:1977:BSS

[385] A. Fielding and C. A. O’Muircheartaigh. Binary segmentation in survey analysis with particular reference to AID. Journal of the Royal Statistical Society. Series D (The Statistician), 26 (1):17–28, March 1977. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988216.

Tam:1977:CSS

[386] S. M. Tam. A case study: Sample design of the Hong Kong 1976 bycensus. Journal of the Royal Statistical Society. Series D (The Statistician), 26 (1):29–39, March 1977. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988217.

Nelder:1977:MDC

[387] J. A. Nelder. Multi-dimensional contingency table with one factor as a response. Journal of the Royal Statistical Society. Series D (The Statistician), 26(1):41–42, March 1977. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988218.

Williams:1977:UMM

[388] Ken Williams. The use of minimal models. Journal of the Royal Statistical Society. Series D (The Statistician), 26(1):43, March 1977. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988219.

Anderson:1977:SCM

[389] O. D. Anderson and M. Williamson.

Some comments on Mark Williamson’s paper “The Biological Interpretation of Time Series Analysis”. Journal of the Royal Statistical Society. Series D (The Statistician), 26(1):46–50, March 1977. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988220.

Anderson:1977:BJA

[390] O. D. Anderson. A Box–Jenkins analysis of the coloured fox data from Nain, Labrador. Journal of the Royal Statistical Society. Series D (The Statistician), 26(1):51–75, March 1977. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988221.

Anderson:1977:BRB

[391] O. D. Anderson. Book review: Mathematical, Statistical and Financial Tables for the Social Sciences, by Z. W. Kmietowicz; Y. Yannoulis. Journal of the Royal Statistical Society. Series D (The Statistician), 26(1):76, March 1977. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988223.

Senga:1977:BRB

[392] W. M. Senga. Book review: Input-Output Analysis and Its Applications, by R. O’Connor; E. W. Henry. Journal of the Royal Statistical Society. Series D (The Statistician), 26(1):76, March 1977. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2988222.

Openshaw:1977:BRB

[393] Stan Openshaw. Book review: Computational Methods of Multivariate Analysis in Physical Geography, by P. M. Mather. Journal of the Royal Statistical Society. Series D (The Statistician), 26(1):76–77, March 1977. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988224.

Anderson:1977:A

[394] Oliver Anderson and I. G. Omuircheartaigh. Announcement. Journal of the Royal Statistical Society. Series D (The Statistician), 26(1):78, March 1977. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2988225.

Anderson:1977:EVS

[395] O. D. Anderson. Editorial: Vale, salve and into mothballs. Journal of the Royal Statistical Society. Series D (The Statistician), 26(1):79–80, March 1977. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988226.

Anonymous:1977:BMa

[396] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 26(1):??, March 1977. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2988227.

Guttman:1977:WWS

[397] Louis Guttman. What is not what in statistics. Journal of the Royal Statistical Society. Series D (The Statistician), 26(2):81–107, June 1977. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987957.

Anonymous:1977:FMb

[398] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 26(2):108, June 1977. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987956.

Kingdon:1977:CTS

[399] J. M. Kingdon. The case for teaching statistics in schools. Journal of the Royal Statistical Society. Series D (The Statistician), 26(2):109–126, June 1977. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987958.

Anderson:1977:IBJ

[400] O. D. Anderson. The interpretation of Box–Jenkins time series models. Journal of the Royal Statistical Society. Series D (The Statistician), 26 (2):127–145, June 1977. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987959.

Hyett:1977:ECP

[401] G. P. Hyett, J. R. McKenzie, and H. P. Wynn. Estimating a change in proportion: Some comments. Journal of the Royal Statistical Society. Series D (The Statistician), 26(2):147–150, June 1977. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987960.

Anonymous:1977:Aa

[402] Anonymous. Announcement. Journal of the Royal Statistical Society. Series D (The Statistician), 26(2):151–152, June 1977. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2987961.

Anonymous:1977:BMb

[403] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 26(2):??, June 1977. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987962.

Cunliffe:1977:OA

[404] Stella V. Cunliffe. Opening address.

Journal of the Royal Statistical Society. Series D (The Statistician), 26(3):153– 154, September 1977. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987894.

Miller:1977:PEW

[405] Ian Miller. The position of the expert witness. Journal of the Royal Statistical Society. Series D (The Statistician), 26 (3):155–162, September 1977. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987895.

Downton:1977:EEW

[406] F. Downton. Experience as an expert witness in gambling cases. Journal of the Royal Statistical Society. Series D (The Statistician), 26(3):163– 179, September 1977. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987896.

Buckland:1977:SLS

[407] W. R. Buckland and W. A. Pridmore.

Statute law and statistical procedures. Journal of the Royal Statistical Society. Series D (The Statistician), 26(3):181– 201, September 1977. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987897.

Anonymous:1977:FMc

[408] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 26(3):202, September 1977. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987893.

Lindley:1977:PL

[409] D. V. Lindley. Probability and the law. Journal of the Royal Statistical Society. Series D (The Statistician), 26(3):203– 220, September 1977. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987898.

Anderson:1977:DPS

[410] O. D. Anderson. On the distributional properties of serial correlations for general time processes. Journal of the Royal Statistical Society. Series D (The Statistician), 26(3):221–230, September 1977. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987899.

Anonymous:1977:BMc

[411] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 26(3):231, September 1977. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987901.

Clarke:1977:BRB

[412] M. R. B. Clarke. Book review: Introduction to Statistics and Computer Programming, by Carl F. Kossack; Claudia I. Henschke. Journal of the Royal Statistical Society. Series D (The Statistician), 26(3):232, September 1977. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987900.

Miller:1977:SCV

[413] J. A. Miller. Statistics — a city view.

Journal of the Royal Statistical Society. Series D (The Statistician), 26(4): 233–242, December 1977. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987809.

Thornton:1977:FSC

[414] M. Thornton. Financial statistics: Collection and analysis. Journal of the Royal Statistical Society. Series D (The Statistician), 26(4):243–257, December 1977. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987810.

Fairclough:1977:PIP

[415] E. H. Fairclough. Personal interviews and postal questionnaires: Some observations and experiences. Journal of the Royal Statistical Society. Series D (The Statistician), 26(4):259– 268, December 1977. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987811.

Artingstall:1977:RTN

[416] R. W. Artingstall. Random thoughts on non-sampling error. Journal of the Royal Statistical Society. Series D (The Statistician), 26(4):269–284, December 1977. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987812.

Anderson:1977:TSA

[417] O. D. Anderson. Time series analysis and forecasting: Another look at the Box–Jenkins approach. Journal of the Royal Statistical Society. Series D (The Statistician), 26(4):285– 303, December 1977. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987813.

Wood:1977:ICO

[418] P. D. P. Wood. On the importance of correct orientation to sex in cattle. Journal of the Royal Statistical Society. Series D (The Statistician), 26(4): 304–306, December 1977. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987814.

Anonymous:1977:BMd

[419] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 26(4):307, December 1977. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987818.

Tweedie:1977:BRB

[420] M. C. K. Tweedie. Book review: Multivariate Analysis in Behavioural Research, by A. E. Maxwell. Journal of the Royal Statistical Society. Series D (The Statistician), 26(4):308, December 1977. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987815.

Fell:1977:BRB

[421] H. A. Fell. Book review: Lady Luck: The Theory of Probability, by W. Weaver. Journal of the Royal Statistical Society. Series D (The Statistician), 26(4): 308–309, December 1977. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987816.

Anonymous:1977:Ab

[422] Anonymous. Announcements. Journal of the Royal Statistical Society. Series D (The Statistician), 26(4):310– 311, December 1977. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987817.

Anonymous:1977:VI

[423] Anonymous. Volume information. Journal of the Royal Statistical Society. Series D (The Statistician), 26(4):312, December 1977. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987807.

Anonymous:1977:FMd

[424] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 26(4):??, December 1977. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987808.

Belson:1978:ICH

[425] William A. Belson. Investigating causal hypotheses concerning delinquent behaviour, with special reference to new strategies in data collection and analysis. Journal of the Royal Statistical Society. Series D (The Statistician), 27 (1):1–25, March 1978. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988249.

Anonymous:1978:FMa

[426] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 27(1):26–54, March 1978. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2988248.

Knight:1978:CFA

[427] Anne Knight. Common factor analysis: Some recent developments in theory and practice. Journal of the Royal Statistical Society. Series D (The Statistician), 27(1):27–42, March 1978. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988250.

Upton:1978:FRM

[428] Graham Upton. Factors and responses in multidimensional contingency tables. Journal of the Royal Statistical Society. Series D (The Statistician), 27 (1):43–48, March 1978. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988251.

Swan:1978:SDC

[429] A. V. Swan. Sample dimensions for the comparison of growth rates (or how long must you watch a rate of change?). Journal of the Royal Statistical Society. Series D (The Statistician), 27 (1):49–53, March 1978. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988252.

Chatfield:1978:ATS

[430] C. Chatfield. On analysing time-series data showing cyclic variation. Journal of the Royal Statistical Society. Series D (The Statistician), 27(1):55–56, March 1978. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2988253.

Cressie:1978:EPD

[431] Noel Cressie. The exponential and power data transformations. Journal of the Royal Statistical Society. Series D (The Statistician), 27(1):57–60, March 1978. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2988254.

Morgan:1978:DPG

[432] D. D. V. Morgan. In defence of Pearson’s goodness of fit statistic. Journal of the Royal Statistical Society. Series D (The Statistician), 27(1):61–62, March 1978. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988255.

Johnson:1978:CLP

[433] N. L. Johnson. A comment on Liddell’s practical test for comparative trials. Journal of the Royal Statistical Society. Series D (The Statistician), 27 (1):63–64, March 1978. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988256.

Anonymous:1978:BMa

[434] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 27(1):??, March 1978. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2988257.

Smith:1978:SAC

[435] T. M. F. Smith. Statistics: The art of conjecture. Journal of the Royal Statistical Society. Series D (The Statistician), 27(2):65–86, June 1978. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987903.

Ehrenberg:1978:GT

[436] A. S. C. Ehrenberg. Graphs or tables. Journal of the Royal Statistical Society. Series D (The Statistician), 27 (2):87–96, June 1978. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987904.

Poloniecki:1978:TAB

[437] J. D. Poloniecki. The two-armed bandit and the controlled clinical trial. Journal of the Royal Statistical Society. Series D (The Statistician), 27 (2):97–102, June 1978. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987905.

Lee:1978:GNA

[438] Sik Yum Lee. The Gauss–Newton algorithm for the weighted least squares factor analysis. Journal of the Royal Statistical Society. Series D (The Statistician), 27(2):103–114, June 1978. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987906.

Abernathy:1978:IPP

[439] James R. Abernathy. The International POPLAB Program: A review. Journal of the Royal Statistical Society. Series D (The Statistician), 27(2):

115–131, June 1978. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987907.

Chatfield:1978:BRB

[440] C. Chatfield. Book review: Elementary Statistics, by Gene R. Sellers. Journal of the Royal Statistical Society. Series D (The Statistician), 27(2):132, June 1978. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987908.

Greenfield:1978:BRBa

[441] Tony Greenfield. Book review: Winning with Statistics: A Painless First Look at Numbers, Ratios, Percentages, Means and Inference, by Richard P. Runyon. Journal of the Royal Statistical Society. Series D (The Statistician), 27(2):132–133, June 1978. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987909.

Porteous:1978:BRB

[442] Hugh L. Porteous. Book review: Further Mathematical Diversions, by Martin Gardner. Journal of the Royal Statistical Society. Series D (The Statistician), 27(2):133–134, June 1978. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987910.

Silvey:1978:BRB

[443] S. D. Silvey. Book review: Probability and Statistical Inference, by R. V. Hogg; E. A. Tanis. Journal of the Royal Statistical Society. Series D (The Statistician), 27(2):134, June 1978. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987911.

Loynes:1978:BRBa

[444] R. M. Loynes. Book review: Basic Probability and Applications, by Miloslav Nosal. Journal of the Royal Statistical Society. Series D (The Statistician), 27 (2):134–135, June 1978. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987912.

Greenfield:1978:BRBb

[445] Tony Greenfield. Book review: Probability and Statistics for Engineers and Scientists, by Ronald E. Walpole; Raymond H. Myers. Journal of the Royal Statistical Society. Series D (The Statistician), 27(2):135–136, June 1978. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987913.

Sprent:1978:BRB

[446] P. Sprent. Book review: Statistics and Probability in Modern Life. Second Edition, by Joseph Newmark. Journal of the Royal Statistical Society. Series D (The Statistician), 27(2):136–137, June 1978. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987914.

Greenfield:1978:BRBc

[447] Tony Greenfield. Book review: Statistics and Research in Urban Administration and Development, by Bernard Benjamin. Journal of the Royal Statistical Society. Series D (The Statistician), 27(2):137, June 1978. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987915.

Harvey:1978:BRB

[448] R. A. Harvey. Book review: Statistical Yearbook 1976 — Member Countries of the Council for Mutual Economic Assistance, by IPC Industrial Press Ltd. Journal of the Royal Statistical Society. Series D (The Statistician), 27(2): 137–138, June 1978. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987916.

Greenfield:1978:BRBd

[449] Tony Greenfield. Book review: Combinatorial Algorithms: Theory and Practice, by Edward M. Reingold; Jurg Nievergeld; Narsingh Deo. Journal of the Royal Statistical Society. Series D (The Statistician), 27(2):138–139, June 1978. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987917.

Dunsmore:1978:BRB

[450] I. R. Dunsmore. Book review: Probability and Statistics, by A. M. Mathai; P. N. Rathie. Journal of the Royal Statistical Society. Series D (The Statistician), 27(2):139, June 1978. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987918.

Anonymous:1978:BMb

[451] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 27(2):140–144, June 1978. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987919.

Anonymous:1978:FMb

[452] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 27(2):??, June 1978. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987902.

Anonymous:1978:E

[453] Anonymous. Editorial. Journal of the Royal Statistical Society. Series D (The Statistician), 27(3/4):145–146, September/December 1978. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988178.

Bartlett:1978:CSA

[454] M. S. Bartlett. Correlation or spectral analysis? Journal of the Royal Statistical Society. Series D (The Statistician), 27(3/4):147–158, September/December 1978. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2988179.

Priestley:1978:NLM

[455] M. B. Priestley. Non-linear models in time series analysis. Journal of the Royal Statistical Society. Series D (The Statistician), 27(3/4):159–176, September/December 1978. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988180.

Plewes:1978:SAU

[456] Thomas J. Plewes. Seasonal adjustment of the U.S. unemployment rate: Introductory remarks. Journal of the Royal Statistical Society. Series D (The Statistician), 27(3/4):177–179, September/December 1978. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988181.

Anonymous:1978:A

[457] Anonymous. Announcement. Journal of the Royal Statistical Society. Series D (The Statistician), 27(3/4):180, September/December 1978. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2988182.

Shiskin:1978:SAU

[458] Julius Shiskin and Thomas J. Plewes.

Seasonal adjustment of the U.S. unemployment rate. Journal of the Royal Statistical Society. Series D (The Statistician), 27(3/4):181–202, September/December 1978. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988183.

Dagum:1978:MFS

[459] Estela Bee Dagum. Modelling, forecasting and seasonally adjusting economic time series with the X-11 ARIMA method. Journal of the Royal Statistical Society. Series D (The Statistician), 27(3/4):203–216, September/December 1978. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2988184.

Akaike:1978:LTS

[460] Hirotugu Akaike. On the likelihood of a time series model. Journal of the Royal Statistical Society. Series D (The Statistician), 27(3/4):217–235, September/December 1978. CODEN ????

ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988185.

Granger:1978:NCT

[461] C. W. J. Granger. New classes of time series models. Journal of the Royal Statistical Society. Series D (The Statistician), 27(3/4):237–253, September/December 1978. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988186.

Anonymous:1978:BMc

[462] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 27(3/4):254–268, September/December 1978. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2988201.

Neave:1978:BRBa

[463] H. R. Neave. Book review: Nonparametric Statistics — A Contemporary Approach, by R. P. Runyon. Journal of the Royal Statistical Society. Series D (The Statistician), 27(3/4):255, September/December 1978. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2988187.

Neave:1978:BRBb

[464] H. R. Neave. Book review: Descriptive Statistics — A Contemporary Approach, by R. P. Runyon. Journal of the Royal Statistical Society. Series D (The Statistician), 27(3/4):256–257, September/December 1978. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2988188.

Clarke:1978:BRB

[465] G. M. Clarke. Book review: Mathematical Statistics: Its Setting and Scope/Univariate Distributions, by Open University Press. Journal of the Royal Statistical Society. Series D (The Statistician), 27(3/4):257, September/December 1978. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988189.

Maher:1978:BRB

[466] Michael J. Maher. Book review: Introductory Statistics: A Decision Map, 2nd edn, by Thad. R. Harshbarger. Journal of the Royal Statistical Society. Series D (The Statistician), 27(3/4):257– 258, September/December 1978. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988190.

Rao:1978:BRB

[467] M. Bhaskara Rao. Book review: Multivariate Analysis — IV, by P. R. Krishnaiah. Journal of the Royal Statistical Society. Series D (The Statistician), 27(3/4):258–259, September/December 1978. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2988191.

Bissell:1978:BRB

[468] Derek Bissell. Book review: Sampling Inspection and Quality Control, 2nd edn, by G. Barrie Wetherill. Journal of the Royal Statistical Society. Series D (The Statistician), 27(3/4):260, September/December 1978. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2988192.

LeRoux:1978:BRB

[469] A. A. Le Roux. Book review: Statistics for Economics: An Introductory Course, by Brinley Davies; John Foad. Journal of the Royal Statistical Society. Series D (The Statistician), 27(3/4):260– 262, September/December 1978. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988193.

Lury:1978:BRB

[470] D. A. Lury. Book review: Studies in the History of Statistics and Probability, Volume II, by Maurice Kendall; R. L. Plackett. Journal of the Royal Statistical Society. Series D (The Statistician), 27(3/4):262–263, September/December 1978. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2988194.

Loynes:1978:BRBb

[471] R. M. Loynes. Book review: Problems of Analytical Statistics, by Yu. V. Linnik. Journal of the Royal Statistical Society. Series D (The Statistician), 27(3/4):263, September/December 1978. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2988195.

Greenfield:1978:BRBe

[472] Tony Greenfield. Book review: Statistics Tables for Mathematicians, Engineers, Economists and the Behavioural and Management Sciences, by H. R. Neave. Journal of the Royal Statistical Society. Series D (The Statistician), 27(3/4):263–264, September/December 1978. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2988196.

Biggins:1978:BRB

[473] J. D. Biggins. Book review: Models of Spatial Processes, by Arthur Getis; Barry Boots. Journal of the Royal Statistical Society. Series D (The Statistician), 27(3/4):264–265, September/December 1978. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988197.

Anderson:1978:BRB

[474] Oliver D. Anderson. Book review: Time Series Analysis: Forecasting and Control, by George E. P. Box; Gwilym M. Jenkins. Journal of the Royal Statistical Society. Series D (The Statistician), 27 (3/4):265, September/December 1978. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988198.

Hook:1978:BRB

[475] G. I. Hook. Book review: Costing Educational Practice, by J. Fielden; P. K. Pearson. Journal of the Royal Statistical Society. Series D (The Statistician), 27(3/4):265–266, September/December 1978. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2988199.

Siday:1978:BRB

[476] S. E. Siday. Book review: Graphical Techniques for Multivariate Data, by Brian Everitt. Journal of the Royal Statistical Society. Series D (The Statistician), 27(3/4):266–267, September/December 1978. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988200.

Anonymous:1978:FMc

[477] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 27(3/4):??, September/December 1978. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2988177.

Bissell:1979:SPM

[478] A. F. Bissell. A semi-parabolic mask for Cusum charts. Journal of the Royal Statistical Society. Series D (The Statistician), 28(1):1–7, March 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2988027.

McClean:1979:ASM

[479] S. I. McClean and D. L. Karageorgos. An age-stratified manpower model applied to the educational system. Journal of the Royal Statistical Society. Series D (The Statistician), 28(1):9–18, March 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2988028.

Kanji:1979:RPS

[480] G. K. Kanji. The role of projects in statistical education. Journal of the Royal Statistical Society. Series D (The Statistician), 28(1):19–27, March 1979. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988029.

Gandy:1979:CCP

[481] R. J. Gandy. The calculation of catchment populations within the National Health Service. Journal of the Royal Statistical Society. Series D (The Statistician), 28(1):29–37, March 1979. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988030. See

comments [541].

Anonymous:1979:FMa

[482] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 28(1):38, March 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2988026.

Harvey:1979:LP

[483] R. A. Harvey. Learning in production. Journal of the Royal Statistical Society. Series D (The Statistician), 28 (1):39–57, March 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988031.

Miller:1979:LE

[484] Ian Miller. Letter to the Editor.

Journal of the Royal Statistical Society. Series D (The Statistician), 28 (1):58, March 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988032.

Porteous:1979:BRBa

[485] H. L. Porteous. Book review: Newer Uses of Mathematics, by J. Lighthill. Journal of the Royal Statistical Society. Series D (The Statistician), 28 (1):59–60, March 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988033.

Huitson:1979:BRBa

[486] A. Huitson. Book review: Basic Statistics for Medical and Social Science Students, by A. E. Maxwell. Journal of the Royal Statistical Society. Series D (The Statistician), 28(1):60, March 1979. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988034.

Anonymous:1979:OPF

[487] Anonymous. Obituary: Professor Frank Sewell Bray. Journal of the Royal Statistical Society. Series D (The Statistician), 28(1):61, March 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2988036.

Huitson:1979:BRBb

[488] A. Huitson. Book review: Maps and Statistics, by Peter Lewis. Journal of the Royal Statistical Society. Series D (The Statistician), 28(1):61, March 1979. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2988035.

Anonymous:1979:BMa

[489] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 28(1):62, March 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2988037.

Anonymous:1979:VIa

[490] Anonymous. Volume information. Journal of the Royal Statistical Society. Series D (The Statistician), 28(1):??, March 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2988025.

Wetherill:1979:ADT

[491] G. B. Wetherill. An analysis of data on trees. Journal of the Royal Statistical Society. Series D (The Statistician), 28(2):63–70, June 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987680.

Greenfield:1979:SIR

[492] A. A. Greenfield. Statisticians in industrial research: The role and training of an industrial consultant. Journal of the Royal Statistical Society. Series D (The Statistician), 28(2):71–82, June 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987681.

Hedges:1979:QWE

[493] B. M. Hedges. Question wording effects: Presenting one or both sides of a case. Journal of the Royal Statistical Society. Series D (The Statistician), 28(2):83–99, June 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987682.

Holden:1979:DUR

[494] K. Holden and D. A. Peel. The determinants of the unemployment rate: Some empirical evidence. Journal of the Royal Statistical Society. Series D (The Statistician), 28(2):101–107, June 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987683.

Craig:1979:NDJ

[495] John Craig. Numeracy and Dr Johnson. Journal of the Royal Statistical Society. Series D (The Statistician), 28 (2):109–117, June 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987684.

Anderson:1979:DBI

[496] O. D. Anderson and J. G. De Gooijer. On discriminating between IMA (1,1) and ARMA (1,1) processes: Some extensions to a paper by Wichern. Journal of the Royal Statistical Society. Series D (The Statistician), 28(2): 119–133, June 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987685.

Butler:1979:SCC

[497] K. L. Butler. Some comments on Cressie’s observations of the exponential and power data transformations. Journal of the Royal Statistical Society. Series D (The Statistician), 28(2): 135–139, June 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987686.

Anonymous:1979:BMb

[498] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 28(2):140–152, June 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987697.

Watts:1979:SRA

[499] Donald G. Watts. Simple revealing approximations to F and χ^{2} percentage point. Journal of the Royal Statistical Society. Series D (The Statistician), 28 (2):141–142, June 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987687.

Brech:1979:OFS

[500] Ronald Brech. Obituary: Frank sewell bray KM, KCSG, FCA, FIS. Journal of the Royal Statistical Society. Series D (The Statistician), 28(2):143–144, June 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987688.

Mangles:1979:BRB

[501] T. H. Mangles. Book review: Statistical Methods for Engineers and Scientists, by A. C. Bajpai; I. M. Calus; J. A. Fairley. Journal of the Royal Statistical Society. Series D (The Statistician), 28(2):145, June 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987689.

Cooper:1979:BRB

[502] I. Cooper. Book review: Programming and Algorithms, by A. J. Guttman. Journal of the Royal Statistical Society. Series D (The Statistician), 28 (2):146, June 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987690.

Hawkes:1979:BRB

[503] Alan G. Hawkes. Book review: Survival Statistics: The Goal and Risk Theory, by Hilary L. Seal. Journal of the Royal Statistical Society. Series D (The Statistician), 28(2):147, June 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2987691.

Maher:1979:BRBa

[504] Michael J. Maher. Book review: Statistics for Technology, by C. Chatfield. Journal of the Royal Statistical Society. Series D (The Statistician), 28(2): 147–148, June 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987692.

Maher:1979:BRBb

[505] Michael J. Maher. Book review: Statistics in Theory and Practice, by L. R. Connor; A. J. H. Morrell. Journal of the Royal Statistical Society. Series D (The Statistician), 28(2):148, June 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987693.

Dunsmore:1979:BRB

[506] R. Dunsmore. Book review: Statistical Analysis: A Decision-Making Approach, by R. Parsons. Journal of the Royal Statistical Society. Series D (The Statistician), 28(2):148–149, June 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987694.

Porteous:1979:BRBb

[507] H. L. Porteous. Book review: Catastrophe Theory and Its Applications, by Tim Poston; Ian Stewart. Journal of the Royal Statistical Society. Series D (The Statistician), 28(2):149–150, June 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987695.

Porteous:1979:BRBc

[508] H. L. Porteous. Book review: Mathematical Carnival, by Martin Gardner. Journal of the Royal Statistical Society. Series D (The Statistician), 28(2): 150–151, June 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987696.

Anonymous:1979:FMb

[509] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 28(2):??, June 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987679.

Pearce:1979:EDR

[510] S. C. Pearce. Experimental design: R.

A. Fisher and some modern rivals. Journal of the Royal Statistical Society. Series D (The Statistician), 28(3):153– 161, September 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987864.

Stoodley:1979:ADT

[511] K. D. C. Stoodley and M. Mirnia. The automatic detection of transients, step changes and slope changes in the monitoring of medical time series. Journal of the Royal Statistical Society. Series D (The Statistician), 28(3):163– 170, September 1979. CODEN ????

ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987865.

Armitage:1979:ADC

[512] P. Armitage. The analysis of data from clinical trials. Journal of the Royal Statistical Society. Series D (The Statistician), 28(3):171–183, September 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987866.

Barnett:1979:FSU

[513] V. Barnett, R. Kay, and P. H. A. Sneath. A familiar statistic in an unfamiliar guise — a problem in clustering. Journal of the Royal Statistical Society. Series D (The Statistician), 28(3): 185–191, September 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987867.

Kenward:1979:IAM

[514] M. G. Kenward. An intuitive approach to the MANOVA test criteria. Journal of the Royal Statistical Society. Series D (The Statistician), 28(3):193– 198, September 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987868.

Haybittle:1979:SCL

[515] J. L. Haybittle and L. S. Freedman. Some comments on the logrank test statistic in clinical trial applications. Journal of the Royal Statistical Society. Series D (The Statistician), 28(3):199– 208, September 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987869.

Idele:1979:LLN

[516] S. I. Idele. The law of large numbers: Some issues of interpretation and application for beginners. Journal of the Royal Statistical Society. Series D (The Statistician), 28(3):209– 217, September 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987870.

Anonymous:1979:BMc

[517] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 28(3):218, September 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987881.

Walters:1979:DAS

[518] D. E. Walters. In defence of the arc sine approximation. Journal of the Royal Statistical Society. Series D (The Statistician), 28(3):219–222, September 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987871.

Rick:1979:BRB

[519] V. E. Rick. Book review: Introduction to Quantitative Economics, by Brian Haines. Journal of the Royal Statistical Society. Series D (The Statistician), 28 (3):223, September 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987872.

Foster:1979:BRB

[520] N. Foster. Book review: SSRC/RSS:

Reviews of United Kingdom Statistical Sources, Vol. 6, by W. F. Maunder; A. B. Atkinson; A. J. Harrison. Journal of the Royal Statistical Society. Series D (The Statistician), 28(3):223– 225, September 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987873.

Bissell:1979:BRBa

[521] Derek Bissell. Book review: Statistical Method for Accounting Students, by D. Pitt Francis. Journal of the Royal Statistical Society. Series D (The Statistician), 28(3):225–226, September 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987874.

Bissell:1979:BRBb

[522] Derek Bissell. Book review: Probability Concepts in Engineering Planning and Design: Vol. I — Basic Principles, by Alfredo H. S. Ang; Wilson H. Tang. Journal of the Royal Statistical Society. Series D (The Statistician), 28 (3):226–227, September 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987875.

Kollerstrom:1979:BRB

[523] Julian Kollerstrom. Book review: Probability, by Raymond K. C. Li. Journal of the Royal Statistical Society. Series D (The Statistician), 28(3):227– 228, September 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987876.

Ubogu:1979:BRB

[524] Roland E. Ubogu. Book review: An Introduction to Quantitative Economics, by Brian Haines. Journal of the Royal Statistical Society. Series D (The Statistician), 28(3):228–229, September 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987877.

Tarling:1979:BRB

[525] Roger Tarling. Book review: Evidence, Proof and Probability, by Richard Eggleston. Journal of the Royal Statistical Society. Series D (The Statistician), 28(3): 229–230, September 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987878.

Biggins:1979:BRB

[526] J. D. Biggins. Book review: Introduction to Probability and Measure, by K. R. Parthasarathy. Journal of the Royal Statistical Society. Series D (The Statistician), 28(3):230, September 1979. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987879.

Greenfield:1979:BRB

[527] Tony Greenfield. Book review: The Art of Problem Solving (Accompanied by Ackoff’s Fables), by Russell L. Ackoff. Journal of the Royal Statistical Society. Series D (The Statistician), 28 (3):230, September 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987880.

Anonymous:1979:FMc

[528] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 28(3):??, September 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2987863.

Daisley:1979:STR

[529] Peter Daisley. Statistical thinking rather than statistical methods. Journal of the Royal Statistical Society. Series D (The Statistician), 28(4):231–239, December 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2988204.

Lenz:1979:STA

[530] H.-J. Lenz. A system-theoretic approach to statistical quality control in industries. Journal of the Royal Statistical Society. Series D (The Statistician), 28(4): 241–252, December 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988205.

Knight:1979:SSA

[531] P. Knight. Statistical sampling in auditing: An Auditor’s viewpoint. Journal of the Royal Statistical Society. Series D (The Statistician), 28(4):253– 266, December 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988206.

Smith:1979:SSA

[532] T. M. F. Smith. Statistical sampling in auditing: A Statistician’s viewpoint. Journal of the Royal Statistical Society. Series D (The Statistician), 28(4): 267–280, December 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2988207.

Moore:1979:SR

[533] P. G. Moore. The shackles of risk.

Journal of the Royal Statistical Society. Series D (The Statistician), 28(4): 281–293, December 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988208.

Anonymous:1979:FMd

[534] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 28(4):294, December 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2988203.

Barback:1979:OSV

[535] R. H. Barback and P. L. Haslett. Official statistics: A view of their value and the “cost” of their provision. Journal of the Royal Statistical Society. Series D (The Statistician), 28(4):295– 307, December 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988209.

Anonymous:1979:N

[536] Anonymous. Notice. Journal of the Royal Statistical Society. Series D (The Statistician), 28(4):308, December 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2988210.

Triggs:1979:BRB

[537] C. M. Triggs. Book review: Problems and Solutions in Theoretical Statistics, by D. R. Cox; D. V. Hinkley. Journal of the Royal Statistical Society. Series D (The Statistician), 28(4):

309, December 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988211.

Plewis:1979:BRB

[538] Ian Plewis. Book review: The Analysis of Cross-Tabulated Data, by Graham J. G. Upton. Journal of the Royal Statistical Society. Series D (The Statistician), 28(4):309–310, December 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2988212.

Anonymous:1979:VIb

[539] Anonymous. Volume information. Journal of the Royal Statistical Society. Series D (The Statistician), 28(4):311– 312, December 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988202.

Anonymous:1979:BMd

[540] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 28(4):??, December 1979. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2988213.

Skidmore:1981:SCC

[541] Steve Skidmore. Some comments on “The Calculation of Catchment Populations within the National Health Service” by Gandy. Journal of the Royal Statistical Society. Series D (The Statistician), 30(1):69–72, March 1981. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987704. See

[481].

Gandy:1981:CCP

[542] R. J. Gandy. The calculation of catchment populations within the National Health Service — Reprise. Journal of the Royal Statistical Society. Series D (The Statistician), 30(4):297– 302, December 1981. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2988067. See [481, 541].